### v3

# Non Null-Controllability Properties of the Grushin-Like Heat Equation on 2D-Manifolds

Abstract
We study the internal non null-controllability properties of the heat equation on 2-dimensional almost-Riemannian manifolds with an interior singularity, and under the assumption that the closure of the control zone does not contain the whole singularity. We show that if locally, around the singularity, the sub-Riemannian metric can be written in a Grushin form, or equivalently the sub-Laplacian writes as a generalized Grushin operator, then, achieving null-controllability requires at least a minimal amount of time. As locally the manifold looks like a rectangular domain, we consequently focus ourselves on the non null-controllability properties of the generalized Grushin-like heat equation on various Euclidean domains.
1 Introduction
Over the past decade, there has been significant attention devoted to exploring the controllability properties of evolution equations associated with degenerate elliptic operators. These are parabolic operators whose symbol can vanish. For the heat equation, this interest has been spurred by positive answers to the question of controllability of the heat equation associated to the Laplace-Beltrami operator on Riemannian manifolds (see e.g. [Fattorini1971-sq, fursikov1996controllability, Lebeau1995-jy, miller2005unique]). A pioneer work for the sub-elliptic heat equation, more precisely the Grushin equation, is very recent (see e.g. [beauchard2014null]), and shows that in this case, at least a minimum amount of time is required. As for the sub-elliptic wave equation on compact manifold, it has been very recently shown in [letrouit2023subelliptic] that it is never null-controllable. However, this question for the sub-elliptic heat in a general context still lacks of answers.
In the present paper, we therefore study the null-controllability properties of the heat equation
$(\partial_{t}-\Delta)f=\mathbf{1}_{\omega}u,$
where $\Delta$ denotes a sub-Laplacian on a two-dimensional almost-Riemannian manifold, under the assumption that, locally, $\Delta$ can be written as a generalized Grushin-type operator.
1.1 The sub-Riemannian setting
We set here the geometric settings of the present paper. For details on sub-Riemannian geometry, see for instance [Jean2014-dz, Agrachev2019-nm].
Let $\mathcal{M}$ be an orientable two-dimensional manifold, with or without boundary, and let $\mathcal{Z}\subset\mathcal{M}$ be a connected embedded one-dimensional submanifold. We will assume throughout this work that $\mathcal{M}$ and $\mathcal{Z}$ satisfy the following.
  1. H0
    Let $\gamma\in\mathbb{N}^{*}$. $\mathcal{M}$ is a complete almost-Riemannian manifold (see [Agrachev2019-nm, Definition 9.1]), of step $1$ on $\mathcal{M}\setminus\mathcal{Z}$, and $\gamma+1$ on $\mathcal{Z}$ (see [Agrachev2019-nm, Definition 3.1]). In particular, $\mathcal{Z}$ is non-characteristic (in the sense of [Franceschi2020-gs, Definition 2.3]). The injectivity radius from $\mathcal{Z}$, that we denote by $\text{inj}(\mathcal{Z})$, is bounded below by some $C>0$, and $\partial\mathcal{Z}\subset\partial\mathcal{M}$.
The condition on the step of the structure implies the Hörmander condition, also known as bracket-generating condition. Namely, the sub-Riemannian structure on $\mathcal{M}$ is locally generated by a family of smooth vector fields such that, at each point, their iterated Lie brackets span the tangent space to $\mathcal{M}$. In our case, we need $\gamma$ iterations of Lie brackets on $\mathcal{Z}$, while none are needed on $\mathcal{M}\setminus\mathcal{Z}$. Given such a family $\mathcal{D}$ of vector fields, we can consider a metric $G$ which is defined to make $\mathcal{D}$ an orthonormal frame on $\mathcal{M}\setminus\mathcal{Z}$. The metric $G$ is Riemannian almost everywhere (except on $\mathcal{Z}$). Thanks to hypothesis H0, Chow-Raschevskii theorem applies, and the notion of (sub-Riemannian) distance associated to the metric $G$ is well-defined on $\mathcal{M}$. We denote it by $d_{\operatorname{sR}}$, and $(M,d_{\operatorname{sR}})$ is a metric space. We also write, for $p\in\mathcal{M}$,
$\delta(p)=\inf\{d_{\operatorname{sR}}(p,q),q\in\mathcal{Z}\}.$
Observe that due to H0, any open connected neighborhood $\mathcal{O}$ of $\mathcal{Z}$ must satisfy $\mathcal{O}\setminus\mathcal{Z}=\mathcal{O}^{-}\sqcup\mathcal{O}^{+}$. We make the following complementary assumption to H0.
  1. H0’
    There exists $R\in(0,\operatorname{inj}(\mathcal{Z})]$ such that, letting $\mathcal{O}=\{p\in\mathcal{M},\delta(p)<R\}$, $\delta$ is smooth in $\mathcal{O}^{\pm}\cup\mathcal{Z}$.
Remark 1.1.
The assumption on the injectivity radius in H0 is automatically verified whenever $\mathcal{Z}$ is compact and without boundary, and $\mathcal{Z}$ is always compact when $\mathcal{M}$ is compact, since $\mathcal{Z}$ is a closed embedded smooth submanifold of $\mathcal{M}$. The smoothness of $\delta$ in H0’ is automatically verified if $\partial\mathcal{M}=\emptyset$ (note that in this case, $\partial\mathcal{Z}=\emptyset)$. In the case with boundary, assumption H0’ is sufficient to ensure that our arguments work under H1H1-glob, but can be relaxed under H1H1-loc (see the assumptions below and Lemma 5.1). Moreover, the assumption $\partial\mathcal{Z}\subset\partial\mathcal{M}$ is to ensure the right geometrical setting under which we can work in the case $\gamma>1$.
The assumption H0 ensures the existence of a double-sided tubular neighborhood around $\mathcal{Z}$, that we denote by $\mathcal{Z}_{L}\subset\{p\in\mathcal{M},\delta(p)<L\}$, for some $0<L<\operatorname{inj}(\mathcal{Z})$, and $\mathcal{Z}_{L}\simeq(-L,L)\times\mathcal{Z}$. Thanks to the smoothness assumption on $\delta$, we actually have that $\mathcal{Z}_{L}=\{p\in\mathcal{M},\delta(p)<L\}$ for $L\in(0,R)$ (see Lemma 5.1). Whenever $\mathcal{Z}$ is compact and without boundary, such a neighborhood always exists under the sole assumption that $\mathcal{Z}$ is non-characteristic, and the set equality $\mathcal{Z}_{L}=\{p\in\mathcal{M},\delta(p)<L\}$ is verified (see e.g. [Franceschi2020-gs, Proposition 3.1]). Whenever $\mathcal{Z}$ is non-compact or has nonempty boundary, its existence is ensured by the assumption on the injectivity radius (see e.g. [rossi2022relative, Theorem 3.7] combined with [Franceschi2020-gs, Proposition 3.1]).
We now introduce $\omega$ to be an open set of $\mathcal{M}$ that satisfies one of the following.
    1. H1-loc
      $\mathcal{M}\setminus\overline{\omega}$ contains a point $p\in\mathcal{Z}$.
    1. H1-glob
      $d_{\operatorname{sR}}(\omega,\mathcal{Z})>0$.
We observe that there exists an open set $\mathcal{U}\subset\mathcal{M}\setminus\overline{\omega}$, that is either
  1. (i)
    a neighborhood of the point $p$ appearing in H1H1-loc,
  2. (ii)
    a neighborhood of $\mathcal{Z}$ if we are under H1H1-glob,
such that $\mathcal{U}$ is diffeomorphic to $(-L,L)\times\Omega_{y}$, where $\Omega_{y}=\mathbb{S}^{1},\ \mathbb{R}$, or a bounded interval (see Figure 1). Indeed, in the setting of H1H1-glob, it is sufficient to choose $\mathcal{U}=\mathcal{Z}_{L}$, with $0<L<\min(d_{\operatorname{sR}}(\omega,\mathcal{Z}),\operatorname{inj}(\mathcal{Z}))$. In the setting of H1H1-loc, we can choose $\mathcal{U}$ to be a neighborhood of such a point $p$, such that it is diffeomorphic to $(-L,L)\times\mathcal{W}$, for some $L>0$, with $\mathcal{W}\subset\mathcal{Z}\setminus\omega$ a relatively compact neighborhood of $p$ in $\mathcal{Z}$. In either case, since $\mathcal{Z}$ is one-dimensional and connected, $\mathcal{U}$ is always diffeomorphic to a set of the form $(-L,L)\times\Omega_{y}$. From now on, the notation $\mathcal{U}$ will designate a tubular neighborhood constructed as above under one of the setting outlined in H1.
(a) An abstract manifold under H1H1-loc.
(b) The Torus under H1H1-glob.
Figure 1: The singularity is in red, the control zone in green, and the tubular neighborhood in blue.
In such a tubular neighborhood, the sub-Riemannian structure is always generated by smooth vector fields of the form $X=\partial_{x}$ and $Y=\tilde{q}(x,y)\partial_{y}$ (see [AgrachevGaussBonnet, Lemma 1, Theorem 1]).
We endow $\mathbb{M}$ with a smooth non-singular measure $\mu$, and we assume that there exists $L_{0}\in(0,\min(d_{\operatorname{sR}}(\omega,\mathcal{Z}),\operatorname{inj}(\mathcal{Z})))$, such that in $\mathcal{U}$ diffeomorphic to $(-L_{0},L_{0})\times\Omega_{y}$, the vector fields $X,Y$, and the measure $\mu$, write as
  1. H2
    $X=\partial_{x}$ and $Y=q(x)r(y)\partial_{y}$, with $q\in C^{\infty}([-L,L])$, $r\in C^{\infty}(\Omega_{y})$, $r>0$ and identically one whenever $\Omega_{y}=\mathbb{R}$, and $q$ satisfying
[…]
  2. H3
    $\mu=h(x)dxdy$, with $h>0$.
Observe that given the form of $Y$ in H2, the assumptions on $q$ and the fact that we need $\gamma$ iteration of Lie brackets only at the singularity $\mathcal{Z}\simeq\{0\}\times\Omega_{y}$ in H0 are equivalent.
We also emphasize that in particular, H2 says that the sub-Riemannian structure on $\mathcal{M}$ is generated, at least locally, by the smooth vector fields $\left\{\partial_{x},q(x)r(y)\partial_{y}\right\}$. Moreover, it implies the assumption on the injectivity radius from $\mathcal{U}\cap\mathcal{Z}$ given in H0.
Remark 1.2.
Under $H1$-glob, our results below will hold true if we write $\mu=h_{1}(x)h_{2}(y)dxdy$, with $h_{2}>0$. For the simplicity of the presentation, we make the choice of taking $\mu$ as prescribed by H3.
We denote by $\Delta$ the sub-Laplacian with respect to $\mu$. That is, for every sufficiently regular function $f$, in $\mathcal{U}$ we have
$\Delta f=\operatorname{div}_{\mu}(\nabla f)=\frac{1}{h(x)}\partial_{x}(h(x)\partial_{x}f)+q(x)^{2}\partial_{y}(r(y)^{2}\partial_{y}f).$ (1)
Set $L^{2}(\mathcal{M}):=L^{2}(\mathcal{M},\mu)$ the set of square integrable, with respect to $\mu$, real-valued functions on $\mathcal{M}$. We consider the Friedrich extension of $\Delta$ with minimal domain $C_{c}^{\infty}(\mathcal{M})$, and we keep this notation. That is, given a generating frame $\{X_{1},...X_{m}\}$ for the sub-Riemannian structure, the form domain of $\Delta$, that is $D\left(\Delta^{1/2}\right)$, is the completion of $C_{c}^{\infty}(\mathcal{M})$ with respect to the horizontal norm
$\|f\|_{D(\Delta^{1/2})}^{2}:=\int_{\mathcal{M}}\sum_{i=1}^{m}|X_{i}f|^{2}\ d\mu,\quad f\in C_{c}^{\infty}(\mathcal{M}).$
The domain of $\Delta$ is $D(\Delta)=\left\{f\in D\left(\Delta^{1/2}\right),\ \Delta f\in L^{2}(\mathcal{M})\right\}$. Similarly, we denote by $\Delta_{\mathcal{U}}$ the restriction of $\Delta$ to $\mathcal{U}$, that is with domain $D(\Delta_{\mathcal{U}})=\left\{f\in D\left(\Delta_{\mathcal{U}}^{1/2}\right),\ \Delta_{\mathcal{U}}f\in L^{2}\left(\mathcal{M}\right)\right\}$.
1.2 The control problem and main result
The goal of this paper is to study the null-controllability properties of the following system. Let $T>0$,
$\left\{\begin{array}[]{lcll}\partial_{t}f-\Delta f&=&u(t,p)\mathbf{1}_{\omega}(p),&(t,p)\in(0,T)\times\mathcal{M},\\ f(t,p)&=&0,&(t,p)\in(0,T)\times\partial\mathcal{M},\text{ if $\partial\mathcal{M}\neq\emptyset,$}\\ f(0,p)&=&f_{0}(p),&p\in\mathcal{M},\end{array}\right$. (5)
where $f_{0}\in L^{2}(\mathcal{M})$, $f\in L^{2}((0,T)\times\mathcal{M})$ is the state, and $u\in L^{2}((0,T)\times\mathcal{M})$ is the control supported in $\omega$.
Remark 1.3.
From H0, the operator $\Delta$ with domain $D(\Delta)$ is densely defined, self-adjoint on $L^{2}(\mathcal{M})$, and generates an analytic semigroup of contractions $(e^{t\Delta})_{t\geq 0}$ on $L^{2}(\mathcal{M})$ ([strichartz, p. 261]). Moreover it is hypoelliptic ([HormanderHypoell, Theorem 1.1]). From [Pazy2012-qf, Sec. 4.2]), system (5) is well posed in the sense that for every $f_{0}\in L^{2}(\mathcal{M})$, there exists a unique solution $f\in C^{0}([0,T],L^{2}(\mathcal{M}))\cap L^{2}((0,T),D(\Delta^{1/2}))$ given by the Duhamel formula.
Definition 1.4 (Null-Controllability).
We say that system (5) is null-controllable in time $T$ from $\omega$ if, for every $f_{0}\in L^{2}(\mathcal{M})$, there exists $u\in L^{2}((0,T)\times\mathcal{M})$ such that the associated solution $f$ of system (5) satisfies $f(T,\cdot,\cdot)=0$ on $\mathcal{M}$.
Let us introduce respectively the minimal time for null-controllability and the Agmon distance as
[…]
$d_{\operatorname{agm}}$ $:x\in(-L,L)\mapsto\int_{\min(0,x)}^{\max(0,x)}|q(s)|\ ds.$ (7)
For a nonempty set $A\subset(-L,L)$, we may write in this paper $d_{\operatorname{agm}}(A)=\inf_{x\in A}d_{\operatorname{agm}}(x)$.
Observe that $T(\omega)=\infty$ if and only if the system of interest is not null-controllable independently of the final time $T>0$.
Theorem 1.5.
Assume H0 and H0’, and that H2-H3 holds in each of the settings of H1. Depending on the value of $\gamma\geq 1$, we have the following statements for system (5).
  1. (i)
    Assume that $\gamma=1$, that we are in the setting of H1H1-loc, and $r$ is identically one. Then, there exists $L\in(0,\min(d_{\operatorname{sR}}(\omega,\mathcal{Z}),\operatorname{inj}(\mathcal{Z}))$ such that
    $T(\omega)\geq\frac{1}{q^{\prime}(0)}\min\{d_{\operatorname{agm}}(-L),d_{\operatorname{agm}}(L)\}.$ (8)
  2. (ii)
    Assume that we are in the setting of $H1$-glob. If $\gamma>1$, then $T(\omega)=\infty$, and if $\gamma=1$, there exists $L\in(0,\min(d_{\operatorname{sR}}(\omega,\mathcal{Z}),\operatorname{inj}(\mathcal{Z}))$ such that
    $T(\omega)\geq\frac{1}{q^{\prime}(0)}\min\{d_{\operatorname{agm}}(-L),d_{\operatorname{agm}}(L)\}.$ (9)
1.3 State of the art
Around the seventies, V.V. Grushin, in [Grušin_1971, Grušin_1970], and M.S. Baouendi, in [BSMF_1967__95__45_0], introduced a class of degenerate hypoelliptic operators $\partial_{x}^{2}-x^{2\gamma}\partial_{y}^{2}$, with $\gamma>0$, now commonly referred to as the classical (Baouendi-)Grushin operators. The first controllability result was given in [beauchard2014null], on the rectangle $(-1,1)\times(0,\pi)$ with Dirichlet boundary conditions. In this study, the control was supported on a vertical strip contained in $(0,1)\times(0,\pi)$, with its closure not intersecting the singularity $\{0\}\times(0,\pi)$. It is showed that when $\gamma<1$, the equation is null-controllable in any time $T>0$, that when $\gamma>1$, null-controllability never occurs, and that when $\gamma=1$, a minimal time is required to have null-controllability. This minimal time, for $\gamma=1$, was latter obtained in [beauchard20152d], with the control supported on two vertical strips, in each side of the singularity, and for which the closure does not intersect the latter. The first result for the classical Grushin equation on $(-1,1)\times\mathbb{S}^{1}$, and on the rectangle, with the control supported on the complement of a horizontal strip, was given in [CRMATH_2017__355_12_1215_0]. It is proved that when $\gamma=1$, null-controllability is never achieved. In this similar setting for the control zone, but this time with the equation posed on $\mathbb{R}^{2}$, the same result was obtained in [LISSYprolate2025]. In [duprez2020control], the minimal time and non null-controllability of the classical Grushin equation were thoroughly explored, considering a broader range of control support configurations.
The Grushin operator was then subsequently generalized by replacing $x^{\gamma}$ by a sufficiently regular function $q$ behaving like $x^{\gamma}$ as $x\rightarrow 0$. Similar results as for the classical operator were obtained. On the rectangle, the problem of boundary null-controllability has been investigated in [beauchard2020minimal]. For the problem of internal null-controllability, for a wide range of considerations for the control zone, results were obtained in [darde2023null]. On the Grushin sphere, endowed with the canonical measure inherited by $\mathbb{R}^{3}$, analogous results were obtained in [tamekue2022null].
Finally, considering the fractional Grushin equation, positive results were obtained very recently on $\mathbb{R}^{n}\times\mathbb{R}^{m}$ or $\mathbb{R}^{n}\times\mathbb{T}^{m}$ in [jaming:hal-04245179] using spectral inequalities, and in [letrouit2023observability] by means of resolvent estimates.
Although the question of controllability properties of the generalized Grushin operator was mainly investigated in some precise settings, the question of giving a geometric interpretation of these results on manifolds remains, as for a geometric interpretation of the null-controllability results of the sub-elliptic heat equation in general.
Concerning other results of null-controllability of a class of degenerate parabolic equation, we can also cite [beauchard2017heat] that study the heat equation on the Heisenberg group, or results concerning the Kolmogorov equation [LEROUSSEAU20163193, beauchardKolmog2015, koenigfractionalheat, beauchardkolmogorov] among others.
1.4 Structure of the paper
In Section 2, we outline the proofs of our main result, and those of our complementary results.
In Section 3, we collect some results concerning the spectral analysis of the operators under consideration. In particular, in Section 3.1.1 and 3.1.2, we respectively remind of the behavior of the first eigenvalue, and exponential decay of the associated first eigenfunction, of the Fourier components of the classical Grushin operator
$\mathsf{G}_{\xi}=-\partial_{x}^{2}+\frac{q^{(\gamma)}(0)^{2}}{(\gamma!)^{2}}\xi^{2}x^{2\gamma},\quad\xi\in\mathbb{R}\setminus\{0\}.$ (10)
In Section 3.2, we provide the asymptotic behavior of a sequence of eigenvalues of the Fourier components of the generalized Grushin operator
$G_{V,\xi}=-\partial_{x}^{2}+\xi^{2}q(x)^{2}+V(x),\quad\xi\in\mathbb{R}\setminus\{0\}.$ (11)
The study of the exponential decays of the associated eigenfunctions are encapsulated within the proofs of the theorems. In Section 3.2.2, we study the behavior of the first eigenvalue of the operator $G_{V,\xi}$ for $\xi\in\mathbb{C}$.
In Section 4, we study the non null-controllability of initial states in form domains, for the generalized Grushin equation posed on Euclidean domains. When the control acts on vertical strips at non-negative distance from the singularity, for $\gamma\geq 1$, we prove a non null-controllability result in Section 4.1. For $\gamma=1$, and the control acting on the complement of a rectangle, we provide a non null-controllability result in Section 4.2.
Finally, in Section 5, we prove Theorem 1.5.
2 Sketch of proofs
Let us begin by outlining the proof of Theorem 1.5, which is presented in Section 5. We fix $\mathcal{U}$ a tubular neighborhood prescribed by H1H1-loc or H1H1-glob, and satisfying additional conditions that will be precised throughout the proof. Theorem 1.5 will be proved by contradiction, together with cutoff arguments.
The underlying idea is the following. We assume that system (5) is null-controllable in time $T>0$ from $\omega$ satisfying one of the conditions of H1. We fix $\mathcal{U}$ as prescribed by the sub-Riemannian setting, and sufficiently small such that at least H2 and H3 hold (and additional conditions made precise in Section 5).
Given an initial state $f_{0}\in L^{2}(\mathcal{U})$, we can construct under the null-controllability assumption of system (5), by means of cutoffs arguments, a control $u$ supported in an arbitrary small neighborhood of $\partial\mathcal{U}\setminus\partial\mathbb{M}$, that we denote by $\tilde{\omega}$, such that the associated solution of
$\left\{\begin{array}[]{lcll}\partial_{t}f-\Delta_{\mathcal{U}}f&=&\mathbf{1}_{\tilde{\omega}}(x,y)u(t,x,y),&(t,x,y)\in(0,T)\times\mathcal{U},\\ f(t,x,y)&=&0,&(t,x,y)\in(0,T)\times\partial\mathcal{U},\\ f(0,x,y)&=&f_{0}(x,y),&(x,y)\in\mathcal{U},\end{array}\right$. (15)
satisfies $f(T)=0$. We call this a reduction process.
[…]
Thus, for the control system (15), with state space $L^{2}(\mathcal{U})$, we study the null-controllability only for sufficiently regular initial states, which in our case, under H1H1-loc, are to be taken in $D\left(\Delta_{\mathcal{U}}^{1/2}\right)$ (see Proposition 5.3). To avoid any ambiguity and to ensure uniformity across the statements and proofs, we will therefore study the null-controllability of initial states in the form domain for any of our control systems in $\mathcal{U}$, regardless of whether we work under H1H1-loc or H1H1-glob.
To study the null-controllability of initial states in the form domain, of system (15), we address the latter within its coordinates representation, where $\Delta_{\mathcal{U}}$ is expressed as (1). After the transformation $f\mapsto\sqrt{h}f$, we are left with addressing the null-controllability properties of initial states taken in the form domain of the spatial operator properly introduced below in (22), of the following system on $L^{2}(\Omega,dxdy)$, where $\Omega=\Omega_{x}\times\Omega_{y}$ is given in the sub-Riemannian setting, the control zone satisfies one of the conditions outlined in H1, and $V\in L^{\infty}(\Omega_{x})$ depends on the measure $h(x)dxdy$. Let $T>0$,
$\left\{\begin{array}[]{lcll}\partial_{t}f-\partial_{x}^{2}f-q(x)^{2}\partial_{y}(r(y)^{2}\partial_{y}f)+V(x)f&=&u(t,x,y)\mathbf{1}_{\tilde{\omega}}(x,y),&(t,x,y)\in(0,T)\times\Omega,\\ f(t,x,y)&=&0,&(t,x,y)\in(0,T)\times\partial\Omega,\\ f(0,x,y)&=&f_{0}(x,y),&(x,y)\in\Omega.\end{array}\right$. (19)
Consequently, Theorem 1.5(i) and 1.5(ii) follow, respectively, from Theorem 2.1 and Theorem 2.2 below.
The reduction process is treated in Section 5.1, while Theorem 1.5(i) and 1.5(ii) are both proved in Section 5.2.
To allow us to work exclusively in the Euclidean setting, let us denote by $G_{V}$ the operator
$G_{V}=-\partial_{x}^{2}-q(x)^{2}\partial_{y}(r(y)^{2}\partial_{y})+V(x),$
with minimal domain $C_{c}^{\infty}(\Omega)$. We consider its Friedrich extension, still denoted by $G_{V}$, and defined by
$\begin{array}[]{ccl}D(G_{V})&=&\left\{f\in D\left(G_{V}^{1/2}\right),\ G_{V}f\in L^{2}(\Omega)\right\}\\[8.0pt] G_{V}&=&-\partial_{x}^{2}f-q(x)^{2}\partial_{y}(r(y)^{2}\partial_{y})+V(x),\end{array}$ (22)
where $D\left(G_{V}^{1/2}\right)$ is the completion of $C_{c}^{\infty}(\Omega)$ with respect to the norm
$\|f\|_{D\left(G_{V}^{1/2}\right)}^{2}=\int_{\Omega}|\partial_{x}f|^{2}+q(x)^{2}r(y)^{2}|\partial_{y}f|^{2}+V(x)|f|^{2}\ dx\ dy,\quad f\in C_{c}^{\infty}(\Omega).$
Theorem 2.1 (Control in the complement of a rectangle).
Let $\gamma=1$. Consider system (19) with $\Omega_{x}=(-L,L)$, $\Omega_{y}=(0,\pi)$, and $\tilde{\omega}$ the complement of a rectangle $a,b]\times\overline{I}$, $a,b>0$, where $\overline{I}$ is a proper closed interval of $\Omega_{y}$. Assume that $q$ satisfies H2, $r$ is identically one on $\Omega_{y}$, and $V\in L^{\infty}(\Omega)$. Then, for any time $T>0$ such that
$T<\frac{1}{q^{\prime}(0)}\min(d_{\operatorname{agm}}(-a),d_{\operatorname{agm}}(b)),$ (23)
[…]
Theorem 2.2 (Control on vertical strips).
Consider system (19) with $\Omega_{x}=(-L,L)$, $\Omega_{y}=(0,\pi)\text{, }\mathbb{S}^{1}$ or $\mathbb{R}$. Assume that $\tilde{\omega}=\tilde{\omega}_{x}\times\Omega_{y}$, with $\operatorname{dist}(\tilde{\omega}_{x},0)>0$, q satisfies H2, and $V\in L^{\infty}(\Omega)$.
  1. (i)
    If $\gamma=1$, then, for any time
    $T<\frac{d_{\operatorname{agm}}(\tilde{\omega}_{x})}{q^{\prime}(0)}$ (24)
[…]
  2. (ii)
[…]
Recall that when $\Omega_{y}=\mathbb{R}$, the function $r$ is assumed to be identically one.
[…]
Remark 2.3.
Theorems 2.1 and 2.2 stay true for initial states in $D(G^{s})$, for any $s\geq 0$, with $D(G^{0})=L^{2}(\Omega)$. This will be clear from the proof for Theorem 2.2. For this statement in the setting of Theorem 2.1, we refer the reader to Remark 4.3.
Remark 2.4.
Smoothness of the functions $q$ and $r$ is not needed in Theorems 2.1 and 2.2. One can simply require $q\in C^{\gamma+1}([-L,L])$ and $r\in C^{1}(\Omega_{y})$.
Remark 2.5.
In the case $V=0$ and for initial states in $L^{2}(\Omega)$, Theorem 2.1 is already known from [darde2023null, Theorem 1.3].
Remark 2.6.
While Theorems 2.1 and 2.2 assume that $\Omega_{x}$ is symmetric with respect to zero, they remain valid when substituting (L,L)(-L,L) with any other interval containing 0 in its interior (see Section 6). Same remark goes for $\Omega_{y}$, that can be replaced by any bounded interval instead of $(0,\pi)$.
As customary, we will adopt the observability perspective. By linearity of our systems, the null-controllability in time $T>0$ of initial states in $D\left(G_{V}^{1/2}\right)$ of system (19) is equivalent to a $D\left(G_{V}^{-1/2}\right)-L^{2}$ observability inequality for its adjoint system (see e.g. [coron2007control, Theorem 2.44]), that we simply call an observability inequality.
Definition 2.7 ($D\left(G_{V}^{-1/2}\right)-L^{2}$ observability).
Let $T>0$. We say that the adjoint system
$\left\{\begin{array}[]{lcll}\partial_{t}f-\partial_{x}^{2}f-q(x)^{2}\partial_{y}(r(y)^{2}\partial_{y}f)+V(x)f&=&0,&(t,x,y)\in(0,T)\times\Omega,\\ f(t,x,y)&=&0,&(t,x,y)\in(0,T)\times\partial\Omega,\\ f(0,x,y)&=&f_{0}(x,y),&(x,y)\in\Omega,\end{array}\right$. (28)
is observable from $\tilde{\omega}$ in time $T$ if there exists $C>0$ such that for every $f_{0}\in L^{2}(\Omega)$, the associated solution satisfies
$\|f(T)\|_{D\left(G_{V}^{-1/2}\right)}^{2}\leq C\int_{0}^{T}\int_{\tilde{\omega}}|f(t,x,y)|^{2}\ dx\ dy\ dt.$ (29)
Let us start with Theorem 2.2 and with $\Omega_{y}$ bounded. Denote by $(\phi_{n})_{n}$ the eigenfunctions of $\partial_{y}(r(y)^{2}\partial_{y})$, associated to $\xi_{n}^{2}\rightarrow+\infty$, which form an Hilbert basis of $L^{2}(\Omega_{y})$. We follow the classical idea, first proposed in [beauchard2014null], of testing the inequality (29) against the sequence of solutions
$g_{n}(t,x,y)=e^{-\lambda_{n}t}v_{n}(x)\phi_{n}(y),$ (30)
where $v_{n}$ is a normalized in norm eigenfunction, associated to $\lambda_{n}$, of
$G_{V,\xi_{n}}=-\partial_{x}^{2}+\xi_{n}^{2}q(x)^{2}+V(x).$ (31)
Assuming that the observability inequality (29) holds from $\tilde{\omega}_{x}\times\Omega_{y}$ in time $T>0$, and testing it against the solutions $g_{n}$ of the form (30), thus implies the existence of a constant $C>0$ such that, for every $n\geq 1$,
$\|g_{n}(T)\|_{D\left(G_{V}^{-1/2}\right)}^{2}=\frac{e^{-2\lambda_{n}T}}{\lambda_{n}}\leq CT\int_{\tilde{\omega}_{x}}|v_{n}(x)|^{2}\ dx\ dt.$ (32)
We therefore have to compare the dissipation on left of (32), with the decay of the eigenfunctions $v_{n}$ on the set $\tilde{\omega}_{x}$ in the limit $n\rightarrow+\infty$. This amounts to a precise spectral analysis of the Fourier components (31) of our generalized Grushin operator, and a precise analysis of the behavior of its eigenfunctions.
To extract the eigenvalue $\lambda_{n}$ and have good estimates on it, the idea is to see the generalized operator $G_{V,\xi_{n}}$ (31) as a perturbation of the classical operator in Fourier $\mathsf{G}_{\xi_{n}}=-\partial_{x}^{2}+\frac{q^{(\gamma)}(0)^{2}}{(\gamma!)^{2}}\xi_{n}^{2}x^{2\gamma}$. Since the classical eigenfunctions concentrate near $x=0$, where the generalized operator behaves like the classical one by assumption H2, we expect $\lambda_{n}$ to behave asymptotically the same way as the sequence of first eigenvalues of $\mathsf{G}_{\xi_{n}}$. The proof culminates with the application of an Agmon-type arguments, showing that the eigenfunctions $v_{n}$ concentrate outside the control zone significantly faster than they decay across the entire domain. This behavior occurs due to the degeneracy of $q$.
The spectral analysis will be performed replacing $\xi_{n}$ by a generic parameter $\xi$. For the Fourier components $\mathsf{G}_{\xi}$ (10) of the classical operator, the analysis is provided in Section 3.1. For the Fourier components $G_{V,\xi}$ (11) of the generalized operator, see Section 3.2. The proof of Theorem 2.2 is provided along with the Agmon-type argument in Section 4.1.
When $\Omega_{y}=\mathbb{R}$, we treat the problem exclusively in Fourier. We choose as solutions of the Fourier transform of system (28) the sequence $g_{n}(t,x,\xi)=v(x,\xi)e^{-\lambda(\xi)t}\psi_{n}(\xi)$, where $v(\cdot,\xi)$ is an eigenfunction of $G_{V,\xi}$ (11), associated to $\lambda(\xi)$, and $\psi_{n}$ is a sequence of cutoffs that localizes in high frequencies. The strategy is then the same as for the case $\Omega_{y}$ bounded.
For Theorem 2.1, we follow [darde2023null]. Much of the work is already carried out in [darde2023null], so we simply show that their strategy still holds in our case. Let us present the idea when $\Omega_{y}=\mathbb{S}^{1}$. By the Fourier decomposition with respect to $y$ presented above, we can choose a sequence of solutions of the form $g_{N}(t,x,y)=\sum_{n\geq N}a_{n}v_{n}(x)e^{iny}e^{-\lambda_{n}t}$, where the sum is finite. Here, $v_{n}$ is chosen to lie within the first eigenspace of the operator $\partial_{x}^{2}+n^{2}q(x)^{2}+V(x)$, associated to the first eigenvalue $\lambda_{n}$ of the latter. Since near $x=0$ we have that $\partial_{x}^{2}+n^{2}q(x)^{2}+V(x)\approx-\partial_{x}^{2}+n^{2}q^{\prime}(0)^{2}x^{2}+V(x)$, in the limit $n\rightarrow+\infty$ classical arguments from complex perturbation theory will show that $\lambda_{n}\sim nq^{\prime}(0)$, as the effect of the potential $V$ becomes more and more negligible in front of $n^{2}q^{2}$. Choosing $v_{n}$ as the spectral projection of $\tilde{v}_{n}(x)=n^{1/4}e^{-nq^{\prime}(0)x^{2}/2}$, we show using an Agmon-type argument that it concentrates near zero as $v_{n}\approx e^{-nd_{\operatorname{agm}}(x)}$ (which can also been show by a standard WKB argument). Hence, for $N$ large enough, our solutions behave like $g_{N}(t,x,y)\approx\sum_{n\geq N}a_{n}e^{-nd_{\operatorname{agm}}(x)}e^{iny}e^{-nq^{\prime}(0)t}$, or more precisely
$g_{N}(t,x,y)=\sum_{n\geq N}a_{n}\gamma_{t,x}(n)e^{n(iy-q^{\prime}(0)t-(1-\varepsilon)d_{\operatorname{agm}}(x))},$
where the $\gamma_{t,x}(n)$ are error terms.
We observe then that after a change of variables, our solutions resemble complex polynomials of the form $\sum_{n\geq N}a_{n}z^{n}$, up to the error terms $\gamma_{t,x}(n)$. The idea is then to show that observability implies an $L^{2}-L^{\infty}$ inequality on these polynomials with a zero of order $N$ at zero, which cannot hold in small times. This implication is carried out up to the possibility of estimating the error term, and is based on rather complicated complex analysis arguments that we shall not exhibit here (see [duprez2020control, darde2023null]). We refer the reader to Section 4.2 (and [duprez2020control, darde2023null]) for more details. The complex spectral analysis, and the proof of Theorem 2.1, are provided respectively in Sections 3.2.2 and 4.2.
3 Spectral analysis
We want to extract a sequence of eigenvalues of the operator $G_{V}=-\partial_{x}^{2}-q(x)^{2}\partial_{y}(r(y)^{2}\partial_{y})+V(x)$ on $L^{2}(\Omega)$, with $\Omega_{y}=(0,\pi)$ or $\mathbb{S}^{1}$, for which we can get sufficiently precise estimates, and for which we can estimate the decay of some associated eigenfunctions in the control zone. We are therefore interested by the eigenvalue problem
$\left\{\begin{array}[]{llll}-\partial_{x}^{2}f-q(x)^{2}\partial_{y}(r(y)^{2}\partial_{y}f)+V(x)f&=&\lambda f,&(x,y)\in\Omega,\\ f(x,y)&=&0,&(x,y)\in\partial\Omega.\end{array}\right$. (35)
As explained in Section 2, this is done by studying the Fourier components of $G_{V}$, that we denoted by $G_{V,\xi_{n}}$. In the analysis of this section, we replace $\xi_{n}$ by a generic parameter $\xi$.
3.1 Preliminaries on the classical Grushin operator
We introduce, for every $\xi\in\mathbb{R}\setminus\{0\}$, and for every $\gamma\geq 1$, the operator $\mathsf{G}_{\xi}$ on $L^{2}(\Omega_{x})$, defined by
$\begin{array}[]{ccl}D(\mathsf{G}_{\xi})&:=&H^{2}\cap H_{0}^{1}(\Omega_{x}),\\ \mathsf{G}_{\xi}u&:=&-u^{\prime\prime}+\frac{q^{(\gamma)}(0)^{2}}{(\gamma!)^{2}}\xi^{2}|x|^{2\gamma}u.\end{array}$ (38)
$\mathsf{G}_{\xi}$ has compact resolvent, is self-adjoint, positive definite, and has discrete spectrum [Berezin1991-nn]. Hence, its first eigenvalue is given by the Rayleigh formula
$\mu_{\xi}=\min\{\mathcal{Q}_{\xi,L}(u,u)\text{; $u\in D(\mathsf{G}_{\xi}),\|u\|_{L^{2}(-L,L)}=1$}\},$ (39)
with
$\mathcal{Q}_{\xi,L}(u,v)=\displaystyle\int_{-L}^{L}u^{\prime}(x)v^{\prime}(x)+\frac{q^{(\gamma)}(0)^{2}}{(\gamma!)^{2}}\xi^{2}|x|^{2\gamma}u(x)v(x)\ dx.$ (40)
3.1.1 Dissipation speed
The following Proposition is already known [beauchard2014null, Proposition 4], but we propose a slightly different proof with a more precise upper bound. This gain of precision will not be of use for our proofs, but we provide it for its own interest. First, we need the following two Lemmas.
Lemma 3.1.
[prandi2018quantum, Lemma 3.5.] Let $\chi$ be a real-valued Lipschitz function with compact support in $\mathbb{R}$. Let $u\in H^{1}(\mathbb{R})$. Then, we have
$\mathcal{Q}_{1,\infty}(\chi u,\chi u)=\mathcal{Q}_{1,\infty}(u,\chi^{2}u)+\langle u,|\chi^{\prime}|^{2}u\rangle_{L^{2}(\mathbb{R})}.$ (41)
Lemma 3.2.
There exists $C>0$, such that for every $R>0$ and $\delta\in(0,1)$, there exist $\chi_{1},\chi_{2}\in C^{\infty}(\mathbb{R})$ two smooth functions such that
  • $0\leq\chi_{i}(x)\leq 1$, for every $x\in\mathbb{R}$, and $i=1,2$,
  • $\chi_{1}=1$ on $\delta R,\delta R]$ and $\chi_{1}=0$ on $\mathbb{R}\setminus[-R,R]$,
  • $\chi_{2}=1$ on $\mathbb{R}\setminus[-R,R]$ and $\chi_{2}=0$ on $\delta R,\delta R]$,
  • $\chi_{1}(x)^{2}+\chi_{2}(x)^{2}=1$, for every $x\in\mathbb{R}$,
  • $\underset{x\in\mathbb{R}}{\sup}|\chi^{\prime}_{i}(x)|\leq C\displaystyle\frac{\pi}{2(1-\delta)R}$, for $i=1,2$.
Proof.
Without loss of generalities, we present the construction for $x>0$. The case $x<0$ follows by symmetry. We first set $f\in C^{\infty}(\mathbb{R})$ to be defined by
$f(x)=\left\{\begin{array}[]{ll}e^{-1/x},&x>0,\\ 0,&x\leq 0.\end{array}\right$. (44)
Next, we define $g\in C^{\infty}(\mathbb{R})$ by
$g(x)=\frac{f(x)}{f(x)+f(1-x)},$ (45)
which satisfies
[…]
Finally, we set
$s(x)=\frac{\pi}{2}g\left(\frac{x-\delta R}{(1-\delta)R}\right),$ (46)
which satisfies $s\in C^{\infty}(\mathbb{R})$ and
[…]
We define $\chi_{1}$ and $\chi_{2}$ on $(0,+\infty)$ by
$\chi_{\xi,1}(x)$ $=\cos(s(x)),$ (47)
$\chi_{\xi,2}(x)$ $=\sin(s(x)).$ (48)
One easily verifies that $\chi_{1}$ and $\chi_{2}$ satisfy the sought properties for $x>0$. ∎
We can now estimate $\mu_{\xi}$.
Proposition 3.3.
There exists $C>0$ such that
$\mu_{\xi}\geq C|\xi|^{\frac{2}{1+\gamma}},\quad\text{for every $|\xi|>0$,}$ (49)
and such that for every $\varepsilon>0$ sufficiently small, there exists $R>0$, such that for every $|\xi|\geq R$, we have
$\mu_{\xi}\leq(1+\varepsilon)(C+\varepsilon)|\xi|^{\frac{2}{1+\gamma}}.$ (50)
Proof.
We start with the lower bound. Let $\tau_{\xi}:=|\xi|^{1/(1+\gamma)}$. Making the change of variable $s=\tau_{\xi}x$, and setting $v(s)=u(\tau_{\xi}^{-1}y)\tau_{\xi}^{-1/2}$, (39) becomes
$\mu_{\xi}=\tau_{\xi}^{2}\min\{\mathcal{Q}_{1,\tau_{\xi}L}(v,v),v\in H^{2}\cap H_{0}^{1}(-\tau_{\xi}L,\tau_{\xi}L),\|v\|_{L^{2}(-\tau_{\xi}L,\tau_{\xi}L)}=1\}.$ (51)
Thus, we have
$\mu_{\xi}\geq\tau_{\xi}^{2}\cdot\overline{\mu},$ (52)
with $\overline{\mu}:=\inf\{\mathcal{Q}_{1,\infty}(v,v),v\in H^{2}(\mathbb{R}),|x|^{\gamma}v\in L^{2}(\mathbb{R}),\|v\|_{L^{2}(\mathbb{R})}=1\}>0$.
We now treat the upper bound. Let $0<\delta<1$. Set $\chi_{\xi,1}$, $\chi_{\xi,2}$ to be the two smooth functions prescribed by Lemma 3.2 with $R=L\tau_{\xi}$. Thus, $\chi_{\xi,1}$ and $\chi_{\xi,2}$ satisfy
  • $0\leq\chi_{\xi,i}(x)\leq 1$, for $i=1,2$,
  • $\chi_{\xi,1}(x)=1$ on $\delta L\tau_{\xi},\delta L\tau_{\xi}]$ and $\chi_{\xi,1}(x)=0$ on $\mathbb{R}\setminus[-L\tau_{\xi},L\tau_{\xi}]$,
  • $\chi_{\xi,2}(x)=1$ on $\mathbb{R}\setminus[-L\tau_{\xi},L\tau_{\xi}]$ and $\chi_{\xi,2}(x)=0$ on $\delta L\tau_{\xi},\delta L\tau_{\xi}]$,
  • $\chi_{\xi,1}(x)^{2}+\chi_{\xi,2}(x)^{2}=1$, for every $x\in\mathbb{R}$ and $|\xi|>0$,
  • $\underset{x\in\mathbb{R}}{\sup}|\chi^{\prime}_{\xi,i}(x)|\leq C\displaystyle\frac{\pi}{2(1-\delta)L\tau_{\xi}}$, for $i=1,2$ and for some $C>0$ independent of $\xi$ and $\delta$.
Let $v\in H^{1}(\mathbb{R})\cap L^{2}(|s|^{2\gamma}ds)$, $\|v\|_{L^{2}(\mathbb{R})}=1$, be the minimizer of $\mathcal{Q}_{1,\infty}(v,v)$. From Lemma 3.1, we derive that
$\mathcal{Q}_{1,\infty}(v,v)=\sum_{i=1}^{2}\mathcal{Q}_{1,\infty}(\chi_{\xi,i}v,\chi_{\xi,i}v)-\sum_{i=1}^{2}\int_{\mathbb{R}}|\chi^{\prime}_{\xi,i}|^{2}|v|^{2}\geq\mathcal{Q}_{1,\infty}(\chi_{\xi,1}v,\chi_{\xi,1}v)-c(\xi)\|v\|_{L^{2}(\mathbb{R})},$ (53)
where $c(\xi):=2\left(C\displaystyle\frac{\pi}{2(1-\delta)L\tau_{\xi}}\right)^{2}$.
Henceforth, by definition of $\mu_{\xi}$, with the minimum taken on all non-zero elements, we have that
$\mu_{\xi}$ $\leq\tau_{\xi}^{2}\frac{\mathcal{Q}_{1,\tau_{\xi}L}(\chi_{\xi,1}v,\chi_{\xi,1}v)}{\|\chi_{\xi,1}v\|_{L^{2}(-\tau_{\xi}L,\tau_{\xi}L)}^{2}}$
$=\tau_{\xi}^{2}\frac{\mathcal{Q}_{1,\infty}(\chi_{\xi,1}v,\chi_{\xi,1}v)}{\|\chi_{\xi,1}v\|_{L^{2}(\mathbb{R})}^{2}}$
[…]
$=\frac{\tau_{\xi}^{2}}{\|\chi_{\xi,1}v\|_{L^{2}(\mathbb{R})}^{2}}\left(\overline{\mu}+c(\xi)\right).$
This proves the upper bound since $\|\chi_{\xi,1}v\|_{L^{2}(\mathbb{R})}\rightarrow 1$ and $c(\xi)\rightarrow 0$ as $|\xi|$ tends to infinity. ∎
In the case $\gamma=1$, we can actually get very much more precise estimates as we know exactly the first eigenfunction of the harmonic oscillator on $\mathbb{R}$. This is the idea of [beauchard2014null, Lemma 4].
Proposition 3.4.
Let $\gamma=1$. Then,
$\mu_{\xi}\sim q^{\prime}(0)|\xi|\text{, as $|\xi|\to+\infty$.}$ (54)
3.1.2 Exponential decay of the eigenfunctions
In this section, we recall the exponential decay for the first eigenfunctions of $\mathsf{G}_{\xi}$, which follows from the above estimates on the first eigenvalues. The propositions in this section are already proved in [beauchard2014null] for $n\in\mathbb{N}^{*}$, $\Omega_{x}=(-1,1)$, and the operator $\partial_{x}^{2}+n^{2}\pi^{2}x^{2\gamma}$, but hold in our case. So we state them, but omit their proofs. We first have the following Lemma, which is proved as [beauchard2014null, Lemma 2].
Lemma 3.5.
There exists a unique non-negative function $v_{\xi}\in L^{2}(\Omega_{x})$ such that $\|v_{\xi}\|_{L^{2}(\Omega_{x})}=1$ and which solves the problem
$\left\{\begin{array}[]{llll}-\partial_{x}^{2}v_{\xi}+\frac{q^{(\gamma)}(0)^{2}}{(\gamma!)^{2}}\xi^{2}|x|^{2\gamma}v_{\xi}&=&\mu_{\xi}v_{\xi},&\text{ if }x\in\Omega_{x},\\ v_{\xi}(\pm L)&=&0,&\text{ if $\Omega_{x}=(-L,L)$,}\\ \lim_{|x|\rightarrow+\infty}v_{\xi}(x)&=&0,&\text{ if }\Omega_{x}=\mathbb{R},\end{array}\right$. (58)
Moreover, $v_{\xi}$ is even.
We now have the following asymptotic $L^{2}(\Omega_{x})$ exponential decay of the $v_{\xi}$’s, as $|\xi|\rightarrow+\infty$, outside any neighborhood of 0 in (L,L)(-L,L). It is proved in [beauchard2014null, Lemma 3].
Proposition 3.6.
For every $|\xi|>0$ large enough, set
$x_{\xi}=\left(\frac{(\gamma!)^{2}\mu_{\xi}}{q^{(\gamma)}(0)^{2}\xi^{2}}\right)^{1/2\gamma}.$ (59)
Then, there exists $R>0$, such that for every $|\xi|>R$, there exists a function $W_{\xi}$ of the form
$W_{\xi}(x)=A_{\xi}e^{-B_{\xi}x^{\gamma+1}},$ (60)
with, for some $C>0$,
$A_{\xi}=2\frac{\sqrt{x_{\xi}}\mu_{\xi}}{(\gamma+1)B_{\xi}x_{\xi}^{\gamma}}e^{B_{\xi}x_{\xi}^{\gamma+1}}\text{ and }B_{\xi}\sim C|\xi|\text{, as $|\xi|\rightarrow+\infty$,}$ (61)
such that for every $x\geq x_{\xi}$,
$v_{\xi}(x)\leq W_{\xi}(x).$ (62)
Moreover, $W_{\xi}\in C^{2}(\{x\geq x_{\xi}\},\mathbb{R})$ is solution of
$\left\{\begin{array}[]{llll}-\partial_{x}^{2}W_{\xi}+\left(\frac{q^{(\gamma)}(0)^{2}}{(\gamma!)^{2}}\xi^{2}|x|^{2\gamma}-\mu_{\xi}\right)W_{\xi}&\geq&0,&\text{ if }x>x_{\xi},\\ W_{\xi}(L)&\geq&0,&\\ \partial_{x}W_{\xi}(x_{\xi})&<&-\sqrt{x_{\xi}}\mu_{\xi}.\end{array}\right$. (66)
Remark 3.7.
Since $v_{\xi}$ is even, we can extend $W_{\xi}$ on $\Omega_{x}\setminus(-x_{\xi},x_{\xi})$ by $W_{\xi}(-x)=W_{\xi}(x)$, and for every $x\in\Omega_{x}\setminus(-x_{\xi},x_{\xi})$, the proposition still holds.
Remark 3.8.
Due to the upper bound (50) on $\mu_{\xi}$, we have that $x_{\xi}\rightarrow 0$. Moreover, due to the behaviour of $\mu_{\xi}$, and $B_{\xi}$, easy computations show that in the limit at infinity for $|\xi|$, we have for some $C_{1},C_{2},C_{3}>0$,
$\begin{array}[]{lcl}x_{\xi}&\sim&C_{1}|\xi|^{-1/(\gamma+1)}\\ A_{\xi}&\sim&C_{2}|\xi|^{\frac{1}{2(\gamma+1)}}\\ B_{\xi}x_{\xi}^{\gamma+1}&\sim&C_{3}\text{, since $B_{\xi}\sim C_{3}|\xi|$}.\end{array}$ (70)
3.2 Dissipation speed for the Fourier components of the generalized operator
We split this section into two parts. Its first part consists in the real spectral analysis of $G_{V,\xi}$, introduced below, which will be of use when the control zone stays at non-negative distance from the singularity in Theorem 2.2. The second part consists in the complex spectral analysis, which will be of use for Theorem 2.1.
We introduce, for a generic complex parameter $\xi\in\mathbb{C}$, such that $\operatorname{Re}(\xi)>0$, and for every $\gamma\geq 1$, the operator $G_{V,\xi}$ on $L^{2}(\Omega_{x})$, defined by
$\begin{array}[]{crl}D(G_{V,\xi})&:=&H^{2}\cap H_{0}^{1}(\Omega_{x}),\\[6.0pt] G_{V,\xi}u&:=&-u^{\prime\prime}+\xi^{2}q(x)^{2}u+V(x)u,\end{array}$ (73)
with $q$ and $V$ satisfying the appropriate assumptions of Theorem 2.1 and 2.2.
We stress that $G_{V,\xi}$ has compact resolvent, and admits an increasing sequence of eigenvalues $\lambda_{k,\xi}\underset{k\rightarrow+\infty}{\longrightarrow}+\infty$ (see [Berezin1991-nn]).
3.2.1 Real spectral analysis
In this section, we assume that $\xi\in\mathbb{R}$. Thus, $G_{V,\xi}$ is self-adjoint and its eigenvalues are real.
Let us recall the following well-known Lemma for self-adjoint operators.
Lemma 3.9.
Let $A$ be a self-adjoint operator with domain $D(A)$, and $\lambda\in\mathbb{R}$. Then
$\operatorname{dist}(\lambda,\sigma(A))\leq\frac{\|(A-\lambda)u\|}{\|u\|},$ (74)
for every $u\in D(A)$.
We can now investigate the behavior of some eigenvalues $\lambda_{\xi}$ of the operator of interest, that will be shown to behave asymptotically like the $\mu_{\xi}$’s introduced in Section 3.1.
Proposition 3.10.
For every $|\xi|>0$ large enough, there exists a constant $C>0$ and an eigenvalue $\lambda_{\xi}$ of $G_{V,\xi}$ that satisfies
$|\lambda_{\xi}-\mu_{\xi}|\leq C|\xi|^{\frac{1}{\gamma+1}}.$ (75)
Namely, for $|\xi|$ large enough, there exists $C_{2}>C_{1}>0$ such that
$C_{1}|\xi|^{2/(\gamma+1)}\leq\lambda_{\xi}\leq C_{2}|\xi|^{2/(\gamma+1)}.$ (76)
Proof.
In the present proof, we shall denote by $C$ any non-negative constant that does not depend on $\xi$, and by $\|\cdot\|$ the $L^{2}(\Omega_{x})$-norm. We denote by $v_{\xi}$ the first eigenfunction, normalized in $L^{2}$-norm, of $\mathsf{G}_{\xi}$. We have, applying $G_{V,\xi}$ to $v_{\xi}$,
$G_{V,\xi}v_{\xi}=\mu_{\xi}v_{\xi}+Qv_{\xi}+Vv_{\xi},$
with
$\begin{array}[]{lcl}Qv_{\xi}&=&\left(q(x)^{2}\xi^{2}-\frac{q^{(\gamma)}(0)^{2}}{(\gamma!)^{2}}\xi^{2}x^{2\gamma}\right)v_{\xi}.\end{array}$
We want to estimate $(G_{V,\xi}-\mu_{\xi})v_{\xi}$ in $L^{2}$-norm, and the Proposition will follow from Lemma 3.9. Let $\varepsilon>0$ sufficiently small. First, we split the integral as
$\|Qv_{\xi}\|^{2}$ $=\int_{\Omega_{x}}|Qv_{\xi}|^{2}\ dx$
$=\int_{\Omega_{x}\setminus(-\varepsilon,\varepsilon)}|Qv_{\xi}|^{2}\ dx+\int_{(-\varepsilon,\varepsilon)}|Qv_{\xi}|^{2}\ dx.$
We remark that by assumption H2 on $q$ near the singularity, all the derivatives of $q^{2}$ up to $2\gamma-1$ are zero at $x=0$. Hence, near zero,
$q(x)^{2}=\frac{q^{(\gamma)}(0)^{2}}{(\gamma!)^{2}}x^{2\gamma}+O\left(|x|^{2\gamma+1}\right).$
Therefore, by normalization and parity of the $v_{\xi}$’s, using the supersolutions of Proposition 3.6, and since $x_{\xi}\rightarrow 0$, for $|\xi|$ large enough we have
$\int_{(-\varepsilon,\varepsilon)}|Qv_{\xi}|^{2}\ dx$ $\leq C|\xi|^{4}\int_{(-\varepsilon,\varepsilon)}|x|^{4\gamma+2}|v_{\xi}|^{2}\ dx$
$=C|\xi|^{4}\left(\int_{0}^{x_{\xi}}|x|^{4\gamma+2}|v_{\xi}|^{2}\ dx+\int_{x_{\xi}}^{\varepsilon}|x|^{4\gamma+2}|v_{\xi}|^{2}\ dx\right)$
$\leq C|\xi|^{4}\left(x_{\xi}^{4\gamma+2}+\int_{x_{\xi}}^{\varepsilon}|x|^{4\gamma+2}W_{\xi}(x)^{2}\ dx\right).$
Using the definition of $W_{\xi}$, and the change of variable $z=x/x_{\xi}$, we get
$\int_{x_{\xi}}^{\varepsilon}|x|^{4\gamma+2}W_{\xi}(x)^{2}\ dx$ $=\int_{x_{\xi}}^{\varepsilon}|x|^{4\gamma+2}A_{\xi}^{2}e^{-2B_{\xi}x^{\gamma+1}}\ dx$
$=\int_{1}^{\varepsilon/x_{\xi}}x_{\xi}^{4\gamma+2}|z|^{4\gamma+2}x_{\xi}A_{\xi}^{2}e^{-2B_{\xi}x_{\xi}^{\gamma+1}z^{\gamma+1}}\ dz$
$\leq x_{\xi}^{4\gamma+2}\int_{1}^{+\infty}|z|^{4\gamma+2}x_{\xi}A_{\xi}^{2}e^{-2B_{\xi}x_{\xi}^{\gamma+1}z^{\gamma+1}}\ dz.$
Now, thanks to Remark 3.8, for $|\xi|$ large enough,
$C\leq B_{\xi}x_{\xi}^{\gamma+1}\text{ and }\ x_{\xi}A_{\xi}^{2}\leq C.$
Hence,
$\int_{x_{\xi}}^{\varepsilon}|x|^{4\gamma+2}A_{\xi}^{2}e^{-2B_{\xi}x^{\gamma+1}}\ dx\leq Cx_{\xi}^{4\gamma+2}\int_{1}^{+\infty}|z|^{4\gamma+2}e^{-Cz^{\gamma+1}}\ dz.$
with the integral on the right-hand side being finite. We therefore have
$\int_{(-\varepsilon,\varepsilon)}|Qv_{\xi}|^{2}\ dx\leq C|\xi|^{4}x_{\xi}^{4\gamma+2}.$ (77)
On the other hand,
$\int_{\Omega_{x}\setminus(-\varepsilon,\varepsilon)}|Qv_{\xi}|^{2}\ dx$ $\leq C|\xi|^{4}\sup_{\Omega_{x}}\left|q(x)^{2}-\frac{q^{(\gamma)}(0)^{2}}{(\gamma!)^{2}}x^{2\gamma}\right|^{2}\int_{\Omega_{x}\setminus(-\varepsilon,\varepsilon)}W_{\xi}(x)^{2}\ dx$
$=C|\xi|^{4}\int_{\varepsilon}^{1}A_{\xi}^{2}e^{-2B_{\xi}x^{\gamma+1}}\ dx$
$\leq C|\xi|^{4}A_{\xi}^{2}e^{-2B_{\xi}\varepsilon^{\gamma+1}}.$
Obviously, by Remark 3.8, the right-hand term decays exponentially fast as $|\xi|\rightarrow+\infty$. Therefore, there exists a constant $C>0$ such that
$\|Qv_{\xi}\|\leq C\xi^{2}x_{\xi}^{2\gamma+1}.$ (78)
Using once again Remark 3.8, we have that
$\xi^{2}x_{\xi}^{2\gamma+1}\leq C\xi^{2-\frac{2\gamma+1}{\gamma+1}}=C\xi^{\frac{1}{\gamma+1}}.$
Finally, we obviously have
$\|Vv_{\xi}\|\leq\sup_{\Omega_{x}}|V(x)|<\infty,$
and (75) follows. Coupling (75) with Proposition 3.3, we directly obtain (76). ∎
3.2.2 Complex spectral analysis
In this subsection, we show that in the case $\gamma=1$, we can obtain more precise estimates on $\lambda_{\xi}$. The results here directly follow from the spectral analysis provided in [darde2023null], in the case $V=0$. As a matter of fact, we show that the spectral analysis from [darde2023null] still holds for our operator.
We first need to reintroduce some objects from [darde2023null], for which we keep as much as possible the same notations. Set I:=(L,L)I:=(-L,L). Denote by $\mathbbm{1}_{I}$ the operator that maps $v\in L^{2}(I)$ to its extension by zero on $\mathbb{R}$, and by $\mathbbm{1}^{*}_{I}$ its adjoint, which maps $v\in L^{2}(\mathbb{R})$ to its restriction on $I$. Set
$\Sigma_{\theta_{0}}:=\{\nu\in\mathbb{C},|\nu|\geq 1,\arg(\nu)\leq\theta_{0}\}\text{, for some $\theta_{0}\in[0,\pi/2)$. }$ (79)
Let $R>0$ and $\varepsilon>0$. For $\nu\in\Sigma_{\theta_{0}}$, set
$\tilde{\mathcal{Z}}_{\nu}=\mathcal{Z}_{q^{\prime}(0)\nu,R,\varepsilon}:=\left\{z\in\mathbb{C},\ |z|\leq Rq^{\prime}(0)|\nu|,\ \text{dist}(z,\sigma(H_{q^{\prime}(0)\nu}))\geq\varepsilon q^{\prime}(0)|\nu|\right\}.$ (80)
We shall write $G_{V,\nu}$, $\nu\in\mathbb{C}$, to be $G_{V,\xi}$ with $\xi$ complex, to keep up with the notations of [darde2023null], and $G_{\nu}$ to be $G_{V,\nu}$ with $V=0$, not to be mistaken with the classical Grushin operator. Also, we denote, for $\beta\in\mathbb{C}$ with $Re(\beta)>0$, the non-self-adjoint harmonic oscillator $H_{\beta}=-\partial_{x}^{2}+\beta^{2}x^{2}$ on $\mathbb{R}$.
Proposition 3.11.
There exists $\nu_{0}\geq 1$ such that for every $\nu\in\Sigma_{\theta_{0}}$, $|\nu|\geq\nu_{0}$ and $z\in\tilde{\mathcal{Z}}_{\nu}$, we have $z\in\rho(G_{V,\nu})$, and
[…]
Proof.
In the present proof, we shall denote by $\|\cdot\|$ the operator norm. From [darde2023null, Proposition 4.1], we know that such a $z$ introduced in the proposition is in fact in $\rho(G_{\nu})$. For such a $z$, we write
$G_{V,\nu}-z=G_{\nu}+V-z=[1+V(G_{\nu}-z)^{-1}](G_{\nu}-z),$
where the potential $V$ can be understood as a bounded linear operator on $L^{2}(I)$. Since we know from [darde2023null, Proposition 4.1] combined with [darde2023null, Proposition C.1 (iv)], that
$\sup_{z\in\tilde{\mathcal{Z}}_{\nu}}\|(G_{\nu}-z)^{-1}\|=\underset{\underset{\nu\in\Sigma_{\theta_{0}}}{|\nu|\rightarrow+\infty}}{O}\left(\frac{1}{|\nu|}\right),$ (82)
it follows that $1+V(G_{\nu}-z)^{-1}$ is an invertible linear operator on $L^{2}(I)$ for $|\nu|$ large enough, and so is $G_{V,\nu}-z$. Moreover,
$\|[1+V(G_{\nu}-z)^{-1}]^{-1}\|=\frac{1}{1-\|V(G_{\nu}-z)^{-1}\|}\rightarrow 1,$
as $|\nu|\rightarrow+\infty$, $\nu\in\Sigma_{\theta_{0}}$, by (82). Hence, since
$(G_{V,\nu}-z)^{-1}=(G_{\nu}-z)^{-1}[1+V(G_{\nu}-z)^{-1}]^{-1},$
it follows that we also have
$\sup_{z\in\tilde{\mathcal{Z}}_{\nu}}\|(G_{V,\nu}-z)^{-1}\|=\underset{\underset{\nu\in\Sigma_{\theta_{0}}}{|\nu|\rightarrow+\infty}}{O}\left(\frac{1}{|\nu|}\right).$ (83)
By triangular inequality, since we already know from [darde2023null, Proposition 4.1] that
[…]
it is sufficient to check that
[…]
but this is a direct consequence of (82) and (83), and the proposition follows. ∎
Now, using basic tools from perturbation theory (see for instance [Kato1995-xh]), or following the exact same proofs of [darde2023null, Proposition 3.6, Proposition 4.2], we get the following corollary.
Corollary 3.12.
In the limit $|\xi|\rightarrow+\infty$, $\xi\in\Sigma_{\theta_{0}}$, we have that $\lambda_{\xi}$ is algebraically simple, and
$\lambda_{\xi}=q^{\prime}(0)\xi+o(|\xi|).$ (84)
4 The Grushin equation on tensorized domains
In this section, we prove the non null-controllability of our generalized Grushin operators for various considerations of the control zone in some Euclidean domains.
4.1 Proof of Theorem 2.2: control on vertical strips
4.1.1 Agmon estimates
Assume that $\Omega_{y}=\mathbb{S}^{1}\text{ or }(0,\pi)$. In this case, we consider a sequence of solutions of the form
$g_{n}(t,x,y)=e^{-\lambda_{n}t}v_{n}(x)\phi_{n}(y),$
for $n$ large, where $\phi_{n}$ is a normalized in norm eigenfunction of $\partial_{y}(r(y)^{2}\partial_{y})$ on $\Omega_{y}$, associated to $\xi_{n}^{2}$, and $v_{n}$ a normalized in norm eigenfunction of $G_{n}:=G_{V,\xi_{n}}$ defined in (73), associated to $\lambda_{n}$ that satisfies the estimate of Proposition 3.10 when $\gamma\geq 1$, and Corollary 3.12 when $\gamma=1$. We abuse the notation here by denoting $\lambda_{n}$ and $v_{n}$ instead of $\lambda_{\xi_{n}}$ and $v_{\xi_{n}}$. We write the control zone as $\omega:=\omega_{x}\times\Omega_{y}$, with $\overline{\omega_{x}}\cap\{x=0\}=\emptyset$.
Introduce the set
$F_{\delta}:=\left\{x\in\Omega_{x},\quad q(x)^{2}\leq\delta\right\}.$ (85)
By assumption H2, there exists $\delta_{0}$, such that for every $0\leq\delta<\delta_{0}$, the set $F_{\delta}$ is a connected neighborhood of 0.
Introduce the Agmon distance between $\omega_{x}$ and $F_{\delta}$, as
$d_{\operatorname{agm},\delta}(\omega_{x},F_{\delta})=\inf_{x\in\omega_{x},y\in F_{\delta}}\int_{\min(x,y)}^{\max(x,y)}(q(s)^{2}-\delta)_{+}^{1/2}\ ds.$ (86)
Observe that if $q$ satisfies H2, then as $\delta\rightarrow 0$ we retrieve the aforementioned Agmon distance (7).
The proof of Theorem 2.2 follows directly from the following proposition for which we follow the ideas of Agmon estimates (see e.g. [Helffer1988-yw, Agmon]).
Proposition 4.1.
There exists $\delta_{0}>0$, such that for every $\delta\in(0,\delta_{0})$, there exist $n_{0}\in\mathbb{N}$ and $C>0$, such that for every $n\geq n_{0}$,
$\int_{\omega_{x}}v_{n}(x)^{2}\ dx\leq Ce^{-2\xi_{n}(1-\delta)d_{\operatorname{agm},\delta}(\omega_{x},F_{\delta})},$ (87)
where $F_{\delta}$ is defined by (85) and satisfies $F_{\delta}\cap\overline{{\omega}_{x}}=\emptyset$, and $d_{\operatorname{agm},\delta}(\omega_{x},F_{\delta})$ is defined by (86).
Proof.
We drop every notation concerning the variable $x$ to simplify the reading. Let $v_{n}=\theta_{n}e^{-\xi_{n}\psi}$ for some sufficiently regular function $\psi$. Then, we have
$v_{n}^{\prime}$ $=\theta_{n}^{\prime}e^{-\xi_{n}\psi}-\xi_{n}\psi^{\prime}\theta_{n}e^{-\xi_{n}\psi},$
$v_{n}^{\prime\prime}$ $=\theta_{n}^{\prime\prime}e^{-\xi_{n}\psi}-2\xi_{n}\theta_{n}^{\prime}\psi^{\prime}e^{-\xi_{n}\psi}-\xi_{n}\psi^{\prime\prime}\theta_{n}e^{-\xi_{n}\psi}+\xi_{n}^{2}\psi^{\prime 2}\theta_{n}e^{-\xi_{n}\psi}.$
Since $(G_{n}-\lambda_{n})v_{n}=0$, we get that $\theta_{n}$ must satisfy
$-\theta_{n}^{\prime\prime}+2\xi_{n}\theta_{n}^{\prime}\psi^{\prime}+(\xi_{n}\psi^{\prime\prime}-\xi_{n}^{2}\psi^{\prime 2}+\xi_{n}^{2}q^{2}+V-\lambda_{n})\theta_{n}=0.$ (88)
Integrating by parts, we first remark that
$\int_{\Omega_{x}}\theta_{n}^{\prime}\psi^{\prime}\theta_{n}=\frac{1}{2}\int_{\Omega_{x}}\frac{d}{dx}(\theta_{n}^{2})\psi^{\prime},$
which implies that
$\int_{\Omega_{x}}\theta_{n}^{\prime}\psi^{\prime}\theta_{n}=-\frac{1}{2}\int_{\Omega_{x}}\psi^{\prime\prime}\theta_{n}^{2}.$
Therefore, multiplying (88) by $\theta_{n}$ and integrating by parts, we get
$\int_{\Omega_{x}}\theta_{n}^{\prime 2}+(\xi_{n}^{2}q^{2}+V-\xi_{n}^{2}\psi^{\prime 2}-\lambda_{n})\theta_{n}^{2}=0,$
which implies, dividing by $\xi_{n}^{2}$,
$\frac{1}{\xi_{n}^{2}}\int_{\Omega_{x}}\theta_{n}^{\prime 2}+\int_{\Omega_{x}}\left(q^{2}-\psi^{\prime 2}+\frac{V}{\xi_{n}^{2}}-\frac{\lambda_{n}}{\xi_{n}^{2}}\right)\theta_{n}^{2}=0,$
and
$\int_{\Omega_{x}}\left(q^{2}-\psi^{\prime 2}+\frac{V}{\xi_{n}^{2}}-\frac{\lambda_{n}}{\xi_{n}^{2}}\right)\theta_{n}^{2}\leq 0.$
Since the only zero for the function $q$ is at $x=0$ by assumption H2, there exists $\delta_{0}>0$ sufficiently small such that for every $\delta\in(0,\delta_{0})$ the set $F_{\delta}:=\{x\in\Omega_{x},q(x)^{2}\leq\delta\}$ is a connected neighborhood of zero, which is decreasing for the inclusion as $\delta\rightarrow 0^{+}$, and moreover satisfies $F_{\delta}\cap\overline{\omega_{x}}=\emptyset$ since $\overline{\omega_{x}}\cap\{0\}=\emptyset$. We choose such a $\delta\in(0,\delta_{0})$.
As we know that $\lambda_{n}/\xi_{n}^{2}$ tends to zero as $n$ tends to infinity thanks to Proposition 3.10, and the fact that $V\in L^{\infty}(\Omega)$, there exists $n_{0}\in\mathbb{N}$ such that for every $n\geq n_{0}$ we have
$\left\{x\in\Omega_{x},q(x)^{2}\leq\frac{\lambda_{n}-V}{\xi_{n}^{2}}\right\}\subset F_{\delta}.$
Thus, for every $n\geq n_{0}$ we get
$\int_{\Omega_{x}}(q^{2}-\psi^{\prime 2}-\delta)\theta_{n}^{2}\leq 0.$
Choose $\psi(x):=(1-\delta)d_{\operatorname{agm},\delta}(x,F_{\delta})$, where $d_{\operatorname{agm}}$ is the degenerated distance induced by the degenerated Agmon metric $(q^{2}-\delta)_{+}dx^{2}$, where $dx^{2}$ is the standard Euclidean metric on $\Omega_{x}$. We note that we have
$|\psi^{\prime}(x)|^{2}\leq(1-\delta)^{2}(q^{2}-\delta)_{+}.$ (89)
Therefore, on the set $F_{\delta}$ we have that $q^{2}-\psi^{\prime 2}-\delta=q^{2}-\delta\leq 0$, and on the complement of $F_{\delta}$ we have
$q^{2}-\psi^{\prime 2}-\delta$ $\geq(1-(1-\delta)^{2})(q^{2}-\delta)_{+}$
$=(2\delta-\delta^{2})(q^{2}-\delta)_{+}$
$\geq 0.$
It follows that
$(2\delta-\delta^{2})\int_{\Omega_{x}\setminus F_{\delta}}(q^{2}-\delta)\theta_{n}^{2}\leq-\int_{F_{\delta}}(q^{2}-\delta)\theta_{n}^{2}.$
Replacing $\theta_{n}$ by its expression $v_{n}(x)e^{\xi_{n}(1-\delta)d_{\operatorname{agm},\delta}(x,F_{\delta})}$, using the fact that $\overline{\omega_{x}}\cap F_{\delta}=\emptyset$, and that $d_{\operatorname{agm},\delta}(x,F_{\delta})=0$ on $F_{\delta}$, we get that on one hand
$-\int_{F_{\delta}}(q^{2}-\delta)\theta_{n}^{2}$ $=-\int_{F_{\delta}}(q^{2}-\delta)v_{n}^{2}$
$\leq\sup_{F_{\delta}}|q^{2}-\delta|,$
since $v_{n}$ is normalized in $L^{2}$-norm. On the other hand,
$(2\delta-\delta^{2})\int_{\Omega_{x}\setminus F_{\delta}}(q^{2}-\delta)\theta_{n}^{2}$ $=(2\delta-\delta^{2})\int_{\Omega_{x}\setminus F_{\delta}}(q^{2}-\delta)v_{n}^{2}e^{2\xi_{n}(1-\delta)d_{\operatorname{agm},\delta}(x,F_{\delta})}$
$\geq(2\delta-\delta^{2})\int_{\omega_{x}}(q^{2}-\delta)v_{n}^{2}e^{2\xi_{n}(1-\delta)d_{\operatorname{agm},\delta}(x,F_{\delta})}$
$\geq(2\delta-\delta^{2})\min_{\omega_{x}}(q^{2}-\delta)e^{2\xi_{n}(1-\delta)d_{\operatorname{agm},\delta}(\omega_{x},F_{\delta})}\int_{\omega_{x}}v_{n}^{2},$
where we stress that $\min_{\omega_{x}}(q^{2}-\delta)>0$. Therefore, we finally get
$\int_{\omega_{x}}v_{n}^{2}\leq\frac{\sup_{F_{\delta}}|q^{2}-\delta|}{(2\delta-\delta^{2})\min_{\omega_{x}}(q^{2}-\delta)}e^{-2\xi_{n}(1-\delta)d_{\operatorname{agm},\delta}(\omega_{x},F_{\delta})},$
which proves the proposition. ∎
4.1.2 Proof of Theorem 2.2
We can now proceed to the proof of Theorem 2.2.
Proof of theorem 2.2 in the case $\Omega_{y}=\mathbb{S}^{1}\text{ or }(0,\pi)$.
Let us split the proof in two parts.
Proof in the case $\gamma>1$
Choose the sequence of solutions $g_{n}$ introduced at the beginning of the section. Assume that system (28) is observable from $\omega$ in time $T>0$. Then, there exists $C>0$ such that for any $n\in\mathbb{N}^{*}$,
$\frac{e^{-2\lambda_{n}T}}{\lambda_{n}}\int_{\Omega_{x}}v_{n}(x)^{2}\ dx$ $\leq C\int_{0}^{T}\int_{\omega_{x}}e^{-2\lambda_{n}t}v_{n}(x)^{2}\ dx$
$\leq CT\int_{\omega_{x}}v_{n}(x)^{2}\ dx.$
Thanks to Proposition 3.10, we can apply Proposition 4.1 that states that there exists two constants $C_{1},C_{2}>0$, such that for every $n$ sufficiently large, since the integral in the left-hand side values one,
$1\leq C_{1}CT\lambda_{n}e^{2\lambda_{n}T}e^{-\xi_{n}C_{2}}.$
Using now the fact that $\lambda_{n}=\mathcal{O}(n^{\frac{2}{\gamma+1}})$ by Proposition 3.10, we have that for some $C_{3}>0$
$1\leq C_{1}C_{3}CTn^{\frac{2}{\gamma+1}}e^{2C_{3}\xi_{n}^{\frac{2}{\gamma+1}}T}e^{-\xi_{n}C_{2}}.$
We therefore simply have to show that for any time $T>0$, the right hand side tends to zero as $n$ tends to infinity. This is straightforward since
$2C_{3}\xi_{n}^{\frac{2}{\gamma+1}}T-\xi_{n}C_{2}=\xi_{n}(\xi_{n}^{\frac{1-\gamma}{\gamma+1}}2C_{3}T-C_{2}),$
becomes negative for any time $T>0$ as $n$ tends to infinity since $\gamma>1$. This disproves the observability inequality for any time $T>0$.
Proof in the case $\gamma=1$.
In this case, the above strategy still works to disprove the observability in small time, and give analogous results as the ones known from [beauchard2014null, darde2023null]. Similarly to the previous proof, assuming that system (28) is observable in time $T>0$ from $\omega$ implies, applying Proposition 4.1, that for $n$ large enough and some constants $C,C_{1}>0$,
$1\leq C_{1}CT\lambda_{n}e^{2\lambda_{n}T}e^{-2\xi_{n}(1-\delta)d_{\operatorname{agm},\delta}(\omega_{x},F_{\delta})},$
with $\delta>0$ an arbitrary small parameter. Thus, for any $T>0$, observability cannot hold if
$e^{2\lambda_{n}T-2\xi_{n}(1-\delta)d_{\operatorname{agm},\delta}(\omega_{x},F_{\delta})}\rightarrow 0,\text{ as $n$ tends to infinity.}$
We use that $\lambda_{n}=q^{\prime}(0)\xi_{n}+o(\xi_{n})$ for large values of $n$ from Corollary 3.12. Letting $\varepsilon>0$, for $n$ large enough we have $\lambda_{n}\leq(1+\varepsilon)\xi_{n}$, and from the above estimate, the observability inequality cannot hold if, for $n$ large,
$2(1+\varepsilon)q^{\prime}(0)\xi_{n}T-2\xi_{n}(1-\delta)d_{\operatorname{agm},\delta}(\omega_{x},F_{\delta})<0.$
That is, we cannot have
$T<\frac{(1-\delta)d_{\operatorname{agm},\delta}(\omega_{x},F_{\delta})}{(1+\varepsilon)q^{\prime}(0)}.$
Since this is true for any $\delta,\varepsilon>0$ arbitrary small, we cannot have
$T<\frac{d_{\operatorname{agm}}(\omega_{x})}{q^{\prime}(0)}.$
This concludes the proof in the case $\gamma=1$. ∎
Let us now proceed to the case of $\Omega_{y}=\mathbb{R}$ and show that it is analogous to the previous case.
Proof of theorem 2.2 in the case $\Omega_{y}=\mathbb{R}$.
Recall that we set $r(y)=1$ on $\mathbb{R}$. Due to the absence of restriction in the $y$-direction for some $x$, the (equivalent) problems of null-controllability and observability are equivalent to their formulation in Fourier components. Namely, (29) is equivalent to
$\|\hat{g}(T)\|_{D\left(G_{V,\xi}^{-1/2}\right)}^{2}\leq C\int_{0}^{T}\int_{\mathbb{R}}\int_{\tilde{\omega}_{x}}|\hat{g}(t,x,\xi)|^{2}\ dx\ d\xi\ dt,$ (90)
with $\hat{g}$ the solution of
$\left\{\begin{array}[]{lcll}\partial_{t}\hat{g}-\partial_{x}^{2}\hat{g}+\xi^{2}q(x)^{2}\hat{g}+V(x)\hat{g}&=&0,&(t,x,\xi)\in(0,T)\times(-L,L)\times\mathbb{R},\\ \hat{g}(t,\pm L,\xi)&=&0,&(t,y)\in(0,T)\times\mathbb{R},\\ \hat{g}(0,x,\xi)&=&\hat{g}_{0}(x,\xi),&(x,\xi)\in(-L,L)\times\mathbb{R},\end{array}\right$. (94)
where $\hat{g}$ is the partial Fourier transform of $g$ with respect to $y$, and $q$ satisfies assumption H2.
As we do not have to come back to the functions in the $y$ variable, we can drop the hat notation. We also write the dependence in $\xi$ as a variable and not as a subscript. By Proposition 3.10, we can extract an eigenvalue of
$G_{V,\xi}=-\partial_{x}^{2}+q(x)^{2}\xi^{2}+V(x),$ (95)
that we denote by $\lambda(\xi)$, and which satisfies for $|\xi|>0$ sufficiently large
$\lambda(\xi)\leq C|\xi|^{2/(\gamma+1)},\quad\text{for some }C>0.$ (96)
In the case where $q$ satisfies H2 with $\gamma=1$, we have from Corollary 3.12 that $\lambda(\xi)$ can be chosen as the first eigenvalue of $G_{V,\xi}$ (95) which satisfies in the limit $\xi\rightarrow+\infty$
$\lambda(\xi)=\xi q^{\prime}(0)+o(|\xi|).$ (97)
We denote by $g(\cdot,\xi)$ that associated eigenfunction normalized in $L^{2}(\Omega_{x})$-norm.
Due to the form of (94), we can localize these eigenfunctions around high frequencies by multiplying them by a smooth compactly supported function. Let $\delta>0$, and set, for $n>0$, $\psi_{n}\in C_{c}^{\infty}(\mathbb{R})$ verifying
$\operatorname{supp}(\psi_{n})=[n-\delta,n+\delta],\qquad 0\leq\psi_{n}\leq 1,\qquad\psi_{n}(x)=1,\ \forall x\in[n-\delta/2,n+\delta/2].$ (98)
Hence, the sequence of functions
$g_{n}(t,x,\xi)=e^{-\lambda(\xi)t}g(x,\xi)\psi_{n}(\xi)$ (99)
are solutions of (94). Plugging these solutions into the observability inequality (90), we get
$\int_{\mathbb{R}^{2}}\frac{e^{-2\lambda(\xi)T}}{\lambda(\xi)}g(x,\xi)^{2}\psi_{n}(\xi)^{2}\ dx\ d\xi\leq C\int_{0}^{T}\int_{\mathbb{R}}\int_{\tilde{\omega}_{x}}e^{-2\lambda(\xi)t}g(x,\xi)^{2}\psi_{n}(\xi)^{2}\ dx\ d\xi\ dt.$
By normalization of $g(\xi,\cdot)$, for $n$ large enough, the left-hand term is bounded below by
$\int_{\mathbb{R}^{2}}\frac{e^{-2\lambda(\xi)T}}{\lambda(\xi)}g(x,\xi)^{2}\psi_{n}(\xi)^{2}\ dx\ d\xi$ $\geq\int_{n-\delta/2}^{n+\delta/2}\int_{\mathbb{R}}\frac{e^{-2\lambda(\xi)T}}{\lambda(\xi)}g(x,\xi)^{2}\ dx\ d\xi$
$=\int_{n-\delta/2}^{n+\delta/2}\frac{e^{-2\lambda(\xi)T}}{\lambda(\xi)}d\xi.$
For the right-hand term, for $n$ large enough, we have
$\int_{0}^{T}\int_{\mathbb{R}}\int_{\tilde{\omega}_{x}}e^{-2\lambda(\xi)t}g(x,\xi)^{2}\psi_{n}(\xi)^{2}\ dx\ d\xi\ dt$ $\leq\int_{0}^{T}\int_{n-\delta}^{n+\delta}\int_{\tilde{\omega}_{x}}e^{-2\lambda(\xi)t}g(x,\xi)^{2}\ dx\ d\xi$
$\leq T\int_{n-\delta}^{n+\delta}\int_{\tilde{\omega}_{x}}g(x,\xi)^{2}\ dx\ d\xi.$
Notice that Proposition 4.1 holds replacing $\xi_{n}$ with a generic parameter $\xi$, for $|\xi|$ large enough. Thus, we can keep bounding the right-hand side above using Proposition 4.1. We will use the same parameter $\delta>0$ arbitrary small in both the definition of $\psi_{n}$ and when applying Proposition 4.1 and Corollary 3.12.
Proof in the case $\gamma>1$.
In this case, we obtain for some constants $C_{0},C_{1},C_{2}>0$ and $n$ sufficiently large, on one hand
$\int_{\mathbb{R}^{2}}\frac{e^{-2\lambda(\xi)T}}{\lambda(\xi)}g(x,\xi)^{2}\psi_{n}(\xi)^{2}\ dx\ d\xi$ $\geq\int_{n-\delta/2}^{n+\delta/2}\frac{e^{-2\lambda(\xi)T}}{\lambda(\xi)}d\xi$
$\geq\frac{e^{-2C_{0}(n+\delta/2)^{2/(\gamma+1)}T}}{\lambda(n+\delta/2)},$
and on the other hand,
$\int_{0}^{T}\int_{\mathbb{R}}\int_{\tilde{\omega}_{x}}e^{-2\lambda(\xi)t}g(x,\xi)^{2}\psi_{n}(\xi)^{2}\ dx\ d\xi\ dt$ $\leq C_{1}T\int_{n-\delta}^{n+\delta}e^{-\xi C_{2}}\ d\xi$
$\leq 2\delta C_{1}Te^{-(n-\delta)C_{2}}.$
Thus, observability cannot hold in any time $T>0$ such that
$2C_{0}(n+\delta/2)^{2/(\gamma+1)}T-(n-\delta)C_{2}<0,\quad\text{as }n\rightarrow+\infty.$ (100)
But since $\gamma>1$, we observe that the above becomes negative in the limit $n\rightarrow+\infty$ for any fixed time $T>0$. Thus, observability never occurs.
Proof in the case $\gamma=1$.
In this case, we have on one hand, for $n$ sufficiently large, by Corollary 3.12,
$\int_{\mathbb{R}^{2}}\frac{e^{-2\lambda(\xi)T}}{\lambda(\xi)}g(x,\xi)^{2}\psi_{n}(\xi)^{2}\ dx\ d\xi$ $\geq\int_{n-\delta/2}^{n+\delta/2}\frac{e^{-2\lambda(\xi)T}}{\lambda(\xi)}d\xi$
$\geq\frac{e^{-2q^{\prime}(0)(1+\delta)(n+\delta/2)T}}{\lambda(n+\delta/2)}.$
On the other hand, by Proposition 4.1,
$\int_{0}^{T}\int_{\mathbb{R}}\int_{\tilde{\omega}_{x}}e^{-2\lambda(\xi)t}g(x,\xi)^{2}\psi_{n}(\xi)^{2}\ dx\ d\xi\ dt$ $\leq C(\delta)T\int_{n-\delta}^{n+\delta}e^{-2\xi(1-\delta)d_{\operatorname{agm}}(\tilde{\omega}_{x},F_{\delta})}\ d\xi$
$\leq 2\delta C(\delta)Te^{-2(n+\delta)(1-\delta)d_{\operatorname{agm}}(\tilde{\omega}_{x},F_{\delta})}.$
Thus, observability cannot hold in any time $T>0$ such that
$\displaystyle 2q^{\prime}(0)(1+\delta)(n+\delta/2)T-2(n+\delta)(1-\delta)d_{\operatorname{agm}}(\tilde{\omega}_{x},F_{\delta})<0,\quad\text{as }n\rightarrow+\infty.$ (101)
It is therefore not hard to see that this cannot hold in any time
$T<\frac{1-\delta}{q^{\prime}(0)(1+\delta)}d_{\operatorname{agm}}(\tilde{\omega}_{x},F_{\delta}).$ (102)
Since $\delta>0$ can be taken arbitrary small, this yields the sought result and concludes the proof. ∎
4.2 Proof of Theorem 2.1: control in the complement of a rectangle
The proof of this theorem is the same as [darde2023null, Theorem 1.3] in the case $V=0$ and for initial states in $L^{2}(\Omega)$, up to some remarks. We therefore only state the modifications to take into account, and omit some details.
We set ourselves on $L^{2}(\Omega,\mathbb{C})$, for which non null-controllability implies the same on $L^{2}(\Omega,\mathbb{R})$. Indeed, the Grushin operator $G_{V}$ having real-valued coefficients, it acts on complex-valued functions as $\hat{G}_{V}(f+ig)=G_{V}f+iG_{V}g$. It inherits all the properties of $G_{V}$, has domain $D(\hat{G}_{V})=D(G_{V})+iD(G_{V})$, and generates a continuous semigroup $(T(t))_{t}$, which satisfies, denoting $(S(t))_{t}$ the one of $G_{V}$, $T(t)(f+ig)=S(t)f+iS(t)g$. Hence, it is not hard to see that observability in time $T>0$ for $G_{V}$ implies the same for $\hat{G}_{V}$. In particular, given a sequence of solutions refuting the observability for $\hat{G}_{V}$, the real-valued functions given by either the real part or imaginary part will disprove the observability on $L^{2}(\Omega,\mathbb{R})$ for $G_{V}$. We shall drop the hat notation for the complexification of $G_{V}$ has there shall not be any confusion.
Proof of Theorem 2.1.
We prove the theorem in the case $\Omega_{y}=\mathbb{S}^{1}$. The case $(0,\pi)$ then follows from [darde2023null, Appendix A]. Given a sequence of complex numbers $(a_{n})_{n}$, a solution of system (19) is given by
$g(t,x,y)=\sum_{n>N}a_{n-1}g(t,x,n)e^{iny},$ (103)
where the sum is taken finite, $N\in\mathbb{N}$ is taken sufficiently large as specified later, and $g(t,x,n)$ is the solution of system
$\left\{\begin{array}[]{lcll}\left(\partial_{t}-\partial_{x}^{2}+n^{2}q(x)^{2}+V(x)\right)g&=&0,&(t,x)\in(0,T)\times(-L,L),\\ g(t,x,n)&=&0,&(t,x)\in(0,T)\times\partial(-L,L),\\ g(0,x,n)&=&g_{0}(x,n),&x\in(-L,L).\end{array}\right$. (107)
Denote by $v_{n}$ an eigenfunction associated to the first eigenvalue $\lambda_{n}$ of the operator $G_{n,V}=-\partial_{x}^{2}+n^{2}q(x)^{2}+V(x)$. Then $g(t,x,n)=e^{-\lambda_{n}t}v_{n}(x)$. More precisely, we choose $v_{n}$ as follows.
Recall that we denote, for $\beta\in\mathbb{C}$ with $Re(\beta)>0$, the non-self-adjoint harmonic oscillator $H_{\beta}=-\partial_{x}^{2}+\beta^{2}x^{2}$ on $\mathbb{R}$. Then $G_{n,V}$ is a perturbation of $H_{q^{\prime}(0)n}$ restricted to (L,L)(-L,L).
Set $\tilde{v}_{n}(x)=n^{1/4}e^{-nq^{\prime}(0)x^{2}/2}$ the first eigenfunction of $H_{q^{\prime}(0)n}$. We choose $v_{n}$ as
$v_{n}:=\Pi_{n}\tilde{v}_{n},$ (108)
where $\Pi_{n}$ is the spectral projector onto the first eigenspace associated to $\lambda_{n}$ of $G_{n,V}$. Our solution $g$ now writes as
$g(t,x,y)=\sum_{n>N}e^{-\lambda_{n}t}a_{n-1}v_{n}(x)e^{iny}.$ (109)
Assume now that observability (29) holds in time $T>0$ from $\tilde{\omega}$.
We plug our solutions (109), with $v_{n}$ defined by (108), into (29), to obtain that there exists $C>0$, such for any solution of the form (109) we have
$\int_{\Omega}|g(T,x,y)|^{2}\ dx\ dy\leq C\int_{0}^{T}\int_{\tilde{\omega}}|g(t,x,y)|^{2}\ dx\ dy\ dt.$ (110)
We will show that (110) implies an $L^{2}-L^{\infty}$ estimates on complex polynomials $p(z)=\sum_{n\geq N}a_{n}z^{n}$ with a zero of order $N$ at zero.
Let us first treat the left-hand side of (110), showing that it can be bounded from below by some $L^{2}$-norm of complex polynomials, following the proof of [darde2023null, Lemma 3.7].
Thanks to Proposition 3.11, it follows from [darde2023null, Proposition 3.6 and Proposition 4.2] that $\|v_{n}\|_{L^{2}(\Omega_{x})}$ is bounded below uniformly in $n$. Thus, thanks to Corollary 3.12 from which we have $\lambda_{n}\sim nq^{\prime}(0)$, and using the uniform lower bound on $\|v_{n}\|_{L^{2}(\Omega_{x})}$, for $\varepsilon>0$, for some constants $c,C_{\varepsilon}>0$, and for $N>0$ sufficiently large, we have
$\|g(T)\|_{D\left(G_{V}^{-1/2}\right)}^{2}\ dx\ dy$ $=\sum_{n>N}\frac{|a_{n-1}|^{2}}{\lambda_{n}}\|v_{n}\|_{L^{2}(\Omega_{x})}^{2}e^{-2\lambda_{n}T}$
$\geq c\sum_{n>N}\frac{|a_{n-1}|^{2}}{\lambda_{n}}e^{-2\lambda_{n}T}$
$\geq ce^{-2C_{\varepsilon}T}\sum_{n>N}\frac{|a_{n-1}|^{2}}{nq^{\prime}(0)(1+\varepsilon)}e^{-2nq^{\prime}(0)(1+\varepsilon)T}$
$\geq\frac{ce^{-2C_{\varepsilon}T}}{q^{\prime}(0)(1+\varepsilon)}\sum_{n>N}\frac{|a_{n-1}|^{2}}{n+1}e^{-2nq^{\prime}(0)(1+\varepsilon)T}$
Therefore, letting $p=\sum_{n\geq N}a_{n}z^{n}$ be a complex polynomial with a zero of order $N$ at zero, since the functions $z\mapsto z^{n}$ are orthogonal in $L^{2}(D(0,R),dm)$ with $dm$ the complex Lebesgue measure, the above computations give
$\int_{\Omega}|g(T,x,y)|^{2}\ dx\ dy\geq C\|p\|^{2}_{L^{2}\left(D\left(0,e^{-q^{\prime}(0)(1+\varepsilon)T}\right)\right)},$ (111)
for some constant $C>0$ that may differ with the one given by the observability inequality.
We now treat the right-hand side of (110), still following the proof of [darde2023null, Lemma 3.7].
Recall that the Agmon distance is defined by (7)
$d_{\operatorname{agm}}:x\in(-L,L)\mapsto\int_{0}^{x}q(s)\ ds.$ (112)
Let $\varepsilon\in(0,1)$, and define $\gamma_{t,x}(n)$ by
$\gamma_{t,x}(n-1):=e^{-t(\lambda_{n}-q^{\prime}(0)n)}v_{n}(x)e^{nd_{\operatorname{agm}}(x)(1-\varepsilon)}.$ (113)
Then, the solutions prescribed by (109) can be written
$g(t,x,y)=\sum_{n>N}a_{n-1}\gamma_{t,x}(n-1)e^{n(iy-q^{\prime}(0)t-(1-\varepsilon)d_{\operatorname{agm}}(x))}$ (114)
Introduce the change of variable
$(x,z)=(x,e^{iy-q^{\prime}(0)t-(1-\varepsilon)d_{\operatorname{agm}}(x)})$ (115)
which is a diffeomorphism from $(0,T)\times\omega$ onto its image that we denote by $\mathcal{P}$, and which satisfies $\mathcal{P}\subset\Omega_{x}\times U$ with (see Figure 2)
$U$ $=D\left(0,e^{-(1-\varepsilon)d_{\operatorname{agm}}(-a)}\right)\cup D\left(0,e^{-(1-\varepsilon)d_{\operatorname{agm}}(b)}\right)\cup\{z\in\mathbb{C},|z|<1,\operatorname{arg}(z)\notin I\}$
$=D\left(0,e^{-(1-\varepsilon)\min(d_{\operatorname{agm}}(-a),d_{\operatorname{agm}}(b))}\right)\cup\{z\in\mathbb{C},|z|<1,\operatorname{arg}(z)\notin I\}.$ (116)
Under the change of variable (115), the right-hand side of (110) becomes, writing $g$ as given in (114),
$\int_{0}^{T}\int_{\omega}|g(t,x,y)|^{2}\ dx\ dy\ dt=\frac{1}{q^{\prime}(0)}\int_{\mathcal{P}}\left|\sum_{n>N}a_{n-1}\gamma_{t,x}(n-1)z^{n}\right|^{2}\frac{1}{|z|^{2}}\ dx\ dm(z).$ (117)
Letting again $p$ be a complex polynomial with a zero of order $N$ at zero, it writes as $p(z)=\sum_{n\geq N}a_{n}z^{n}$. We see that the right-hand side above is almost an integral on $p$, and we will try to bound it from above by some $L^{\infty}$-norm of $p$.
Introduce the operator $\gamma_{t,x}$ acting on complex polynomials as
$\gamma_{t,x}(z\partial_{z})\left(\sum_{n}a_{n}z^{n}\right)=\sum_{n}a_{n}\gamma_{t,x}(n)z^{n}.$ (118)
This operator is introduced in [darde2023null, eq. (3.19)] in the case $V=0$, and its properties are given by [darde2023null, Lemma 3.5]. Therefore, (117) is bounded as
$\int_{0}^{T}\int_{\tilde{\omega}}|g(t,x,y)|^{2}\ dx\ dy\ dt$ $=\frac{1}{q^{\prime}(0)}\int_{\mathcal{P}}\left|\gamma_{t,x}(z\partial_{z})(p)(z)\right|^{2}\ dx\ dm(z)$
$\leq C\sup_{(t,x)\in(0,T)\times\Omega_{x}}\|\gamma_{t,x}(z\partial_{z})(p)\|_{L^{\infty}(U)}.$
We have now arrived to the important point of the proof. We need to bound one last time the right-hand side above by
$\sup_{(t,x)\in(0,T)\times\Omega_{x}}\|\gamma_{t,x}(z\partial_{z})(p)\|_{L^{\infty}(U)}\leq C\|p\|_{L^{\infty}(\mathcal{V})},$ (119)
for some constant $C>0$ independent of $(t,x)$, and $\mathcal{V}$ an arbitrary small neighborhood of $\overline{U}$.
Let us admit for now that this is possible and conclude the proof of Theorem 2.1, following the exact same lines as for the proof of [darde2023null, Theorem 1.3, Theorem 3.1]. We have seen that if observability (29) holds in time $T>0$ from $\tilde{\omega}$, then from (111) and the admitted upper bound (119), there exists $C>0$ such that for every complex polynomial $p$ with a zero of order $N$ at zero, $N$ large enough, we have
$\|p\|^{2}_{L^{2}\left(D\left(0,e^{-q^{\prime}(0)(1+\varepsilon)T}\right)\right)}\leq C\|p\|_{L^{\infty}(\mathcal{V})},$ (120)
where we recall that $\mathcal{V}$ is an arbitrary small neighborhood of $\overline{U}$ defined in (4.2). For every $\varepsilon>0$, for every $T>0$ such that
$T\leq\frac{1-\varepsilon}{q^{\prime}(0)(1+\varepsilon)}\min(d_{\operatorname{agm}}(-a),d_{\operatorname{agm}}(b)),$ (121)
there exists $z_{0}\in D(0,e^{-q^{\prime}(0)(1+\varepsilon)T})\setminus\overline{U}$ (see Figure 2). Hence, by Runge’s theorem, there exists a sequence of polynomials $(\tilde{p}_{k})_{k}$ that converges to $1/(z-z^{0})$, uniformly on every compact subset of $\mathbb{C}\setminus z_{0}[1,+\infty)$. We then choose the sequence of polynomials $p_{k}=z^{N+1}\tilde{p}_{k}$ to disprove (120). Indeed, as $k\rightarrow+\infty$, the left-hand side of (120) explodes, while the right-hand side remains uniformly bounded.
Now, since $\varepsilon$ can be chosen arbitrary small, the result of Theorem 2.1 follows.
Figure 2: The yellow pacman shape is $U$. The disk in red is $D\left(0,e^{-(1-\varepsilon)\min(d_{\operatorname{agm}}(-a),d_{\operatorname{agm}}(b))}\right)$, and it contains the dotted disk $D\left(0,e^{-q^{\prime}(0)(1+\varepsilon)T}\right)$. When $T$ is too small, there exists $z_{0}\in D(0,e^{-q^{\prime}(0)(1+\varepsilon)T})\setminus\overline{U}$.
In particular, the sequence of complex numbers used in the solutions of the form (109) to disprove the observability inequality (110) are entirely determined by the sequence of polynomials $(\tilde{p}_{k})_{k}$ above.
To fully conclude on the present proof, recall that we are left with proving that for any complex polynomial with a zero of order $N$ at zero, with $N$ sufficiently large, we have
$\sup_{(t,x)\in(0,T)\times\Omega_{x}}\|\gamma_{t,x}(z\partial_{z})(p)\|_{L^{\infty}(U)}\leq C\|p\|_{L^{\infty}(\mathcal{V})}.$ (122)
This is [darde2023null, Lemma 3.5] which follows from [darde2023null, Theorem 4.7], proved in [CRMATH_2017__355_12_1215_0, Theorem 18], and the rest of this proof is dedicated to proving (122).
The idea is to observe that, at least not requiring for now the independence of $C>0$ on $(t,x)$ for now, this is possible if the function
$z\mapsto\sum_{n\geq N}\gamma_{t,x}(n)z^{n}$
admits an holomorphic extension on $\mathbb{C}\setminus[1,+\infty)$. It has been shown in [CRMATH_2017__355_12_1215_0, Theorem 18] that a sufficient condition is that the error term $\gamma_{t,x}$ can be interpolated by some function in a certain class of holomorphic functions $\mathcal{S}(D)$ that we introduce below.
For the constant $C$ to be independent of $t,x$, we need the family of interpolations to be uniformly bounded in $\mathcal{S}(D)$ with respect to $t$ and $x$.
From Section 3.2.2, for every $\theta\in[0,\frac{\pi}{2})$ there exists $r_{\theta}\geq 1$ such that Proposition 3.11 holds for any $\nu\in\Sigma_{\theta_{0}}\setminus D(0,r_{\theta_{0}})$. Following [darde2023null, Definition 4.6], we define
$D:=\cup_{\theta\in[0,\pi/2)}\Sigma_{\theta}\setminus D(0,r_{\theta}),$
and $\mathcal{S}(D)$ to be the set of holomorphic function $f$ on $D$ such that for every $\theta\in[0,\pi/2),\delta>0$ we have
$\sup_{D\cap\Sigma_{\theta}}\left|f(z)e^{-\delta|z|}\right|<+\infty.$
Observe that $\mathcal{S}(D)$ is stable under multiplication. Hence, it is sufficient to show separately that both terms $e^{-t(\lambda_{n}-q^{\prime}(0)n)}$ and $v_{n}(x)e^{nd_{\operatorname{agm}}(x)(1-\varepsilon)}$ forming $\gamma_{t,x}(n)$ (113) can be interpolated by an element of $\mathcal{S}(D)$.
Let us treat the first term $e^{-t(\lambda_{n}-q^{\prime}(0)n)}$. Write $\lambda(z)$ to be the first eigenvalue of $\partial_{x}^{2}+z^{2}q(x)^{2}+V(x)$ obtained in Section 3.2.2. With this notation, we have $\lambda_{n}=\lambda(n)$. Hence, the first term is interpolated by the function
$\mathcal{E}_{t}:z\mapsto e^{-t(\lambda(z)-q^{\prime}(0)z)}.$ (123)
Similarly, denoting $v_{z}=\Pi_{z}\left(z^{1/4}e^{-zq^{\prime}(0)x^{2}/2}\right)$, the second term is interpolated by
$\mathcal{F}_{x}:z\mapsto v_{z}(x)e^{zd_{\operatorname{agm}}(x)(1-\varepsilon)}.$ (124)
From complex perturbation theory (see e.g. [Kato1995-xh, Chapter 7.1]), both functions $\mathcal{E}_{t}$ and $\mathcal{F}_{x}$ are holomorphic.
Using Proposition 3.11, we see that the interpolations $\mathcal{E}_{t}$ belong to $\mathcal{S}(D)$ and are uniformly bounded. Similarly, thanks to Proposition 4.2 below, the interpolations $\mathcal{F}_{x}$ belong to $\mathcal{S}(D)$ and are uniformly bounded. For details on these last two facts, we refer the reader to [darde2023null, Lemma 3.5]. Thus, estimate (122) is true, which concludes the proof.
Proposition 4.2.
Let $\theta_{0}\in[0,\pi/2)$, and $\varepsilon\in(0,\pi)$. There exists $C>0$ such that for $\nu\in\Sigma_{\theta_{0}}$, $|\nu|$ large enough, we have
$\int_{I}\left|v_{\nu}(x)e^{\nu(1-\varepsilon)d_{\operatorname{agm}}(x)}\right|^{2}\ dx$ $\leq C|\nu|,$ (125)
$\|v_{\nu}(x)e^{\nu(1-\varepsilon)d_{\operatorname{agm}}(x)}\|_{L^{\infty}(I)}$ $\leq C|\nu|.$ (126)
Proof.
We have the following Agmon equality from the exact same computations as in [darde2023null, Proposition 4.3],
$\|v_{\nu}^{\prime}\|_{L^{2}(I)}^{2}+\int_{I}(|\nu|^{2}q(x)(1-(1-\varepsilon)^{2})+V-\mu_{\nu})|v_{\nu}(x)|^{2}\ dx=0,$
with
$\mu_{\nu}:=\frac{Re(e^{-i\arg(\nu)}\lambda_{\nu})}{\cos(\arg(\nu))}.$
For some constant $C>0$, for $|\nu|$ large enough, we have for some $c>0$ and every $x$ such that $d_{\operatorname{agm}(x)}|\nu|>C$,
$|\nu|^{2}q(x)(1-(1-\varepsilon)^{2})+V-\mu_{\nu}\geq c|\nu|,$
since in the limit, the potential, whether positive or negative, becomes negligible in front of the other terms. The proposition follows now from the exact same arguments as in [darde2023null, Corollary 4.5]. ∎
Remark 4.3.
We observe from the above proof that Theorem 2.1 is still true if we consider initial states in $D(G^{s})$, for any $s>0$. In this case, we have in the left-hand side of the observability inequality (29) a $D(G^{-s})$-norm, and still an $L^{2}((0,T)\times\omega)$ norm in the right-hand side. Thus, we just have to bound from below the left-hand side. We have, following the computations of the beginning of the proof,
$\|g(T)\|_{D(G^{-s})}^{2}\ dx\ dy$ $=\sum_{n>N}\frac{|a_{n-1}|^{2}}{{\lambda_{n}}^{2s}}\|v_{n}\|_{L^{2}(\Omega_{x})}^{2}e^{-2\lambda_{n}T}$
$\geq\frac{Ce^{-2C_{\varepsilon}T}}{q^{\prime}(0)^{2s}(1+\varepsilon)^{2s}}\sum_{n>N}\frac{|a_{n-1}|^{2}}{n^{2s}+1}e^{-2nq^{\prime}(0)(1+\varepsilon)T}$
$=\frac{Ce^{-2C_{\varepsilon}T}}{q^{\prime}(0)^{2s}(1+\varepsilon)^{2s}}\sum_{n\geq N}\frac{|a_{n}|^{2}}{(n+1)^{2s}+1}e^{-2(n+1)q^{\prime}(0)(1+\varepsilon)T}.$
Let $T>0$ be an horizon time for which we know that (120) cannot hold. Denote by $(p_{k})_{k}$ the sequence of polynomial refuting the inequality (120) in the above proof. That is, $p_{k}=z^{N+1}\tilde{p}_{k}=z^{N+1}\sum_{n}a_{n,k}z^{n}$, where $\tilde{p}_{k}$ converges uniformly on every compact subsets of $\mathbb{C}\setminus(z_{0}[1,+\infty))$ to $(z-z_{0})^{-1}$, with $z_{0}$ chosen as in the proof of Theorem 2.1. We stress that in particular, we have $|z_{0}|<e^{-q^{\prime}(0)(1+\varepsilon)T}$. Thus, the solutions constructed from the proof of Theorem 2.1 that refutes the observability (29) are given by
$g_{k}(t,x,y)=\sum_{n\geq N}a_{n-N,k}v_{n+1}(x)e^{-\lambda_{n+1}t+i(n+1)y}.$
Let us see that these $g_{k}$ also disprove the $D(G^{-s})-L^{2}$ observability. If observability would hold in time $T>0$, we would obtain that there exists $C>0$, different from the observability constant, such that
$\sum_{n>N}\frac{|a_{n-N-1,k}|^{2}}{n^{2s}+1}e^{-2nq^{\prime}(0)(1+\varepsilon)T}\leq C\int_{0}^{T}\int_{\omega}|g_{k}(t,x,y)|^{2}\ dx\ dy\ dt,$ (127)
with the right-hand side uniformly bounded. We need to show that the left-hand side diverges as $k\rightarrow+\infty$. By the Cauchy formula, and by the uniform convergence of $(\tilde{p}_{k})$ to $(z-z_{0})^{-1}$, we have that for any $n\geq 0$, $a_{n,k}$ converges to $1/z_{0}^{n+1}$. Thus,
$\frac{|a_{n-N-1,k}|^{2}}{n^{2s}+1}\rightarrow\frac{1}{n^{2s}}\left|\frac{1}{z_{0}}\right|^{2(n-N)},\quad\text{as}\quad k\to+\infty.$
By Fatou’s Lemma, we see that the left-hand term in (127) will diverge to infinity if $|z_{0}|<e^{-q^{\prime}(0)(1+\varepsilon)T}$, which is the case by our choice of horizon time $T>0$. This concludes.
Remark 4.4.
It would be interesting to obtain this result when $r$ is not identically one, which seems to be a complicated task. Indeed, the holomorphic interpolation argument becomes a challenging issue. In this case, the first eigenvalue $\lambda_{n}$ becomes $\lambda_{\xi_{n}}$, where we recall that $\xi_{n}^{2}$ is the $n$-th eigenvalue of $\partial_{y}(r(y)^{2}\partial_{y})$, and the associated eigenfunctions are now perturbations of sinusoidal, at least for large values of $n$. Ignoring the eigenfunctions issue as we believe it is not the most important one, for almost every $(t,x)$ we want to find an holomorphic function $F_{t,x}\in\mathcal{S}(D)$ such that $F_{t,x}(n)=\gamma_{t,x}(n)$. This first amounts to interpolate for example the term $e^{-t(\lambda_{\xi_{n}}-q^{\prime}(0)n)}$ at each $n$. We saw that when $r=1$, the interpolation is directly given by the same formula replacing $n$ by $z$, as the holomorphy is ensured by perturbation theory arguments. But when $r$ is not identically one, we must find $G$, holomorphic in the right half-plane (see [CRMATH_2017__355_12_1215_0]), such that $G(n)=\xi_{n}$. Indeed, in this case, the composite function $z\mapsto G(z)\mapsto\lambda(G(z))$ interpolates correctly $\lambda_{\xi_{n}}$. The existence of an interpolation is not a complicated problem, but one must keep in mind that we have heavy restrictions on it since the interpolation of the error term must in the end belong to $\mathcal{S}(D)$.
5 Reduction to a local problem and proof of Theorem 1.5
In this section, we prove our main result, Theorem 1.5. Throughout this section, we set ourselves under the geometric setting prescribed by Section 1.1. Recall that, according to the statement of Theorem 1.5, when working under H1H1-loc, our almost-Riemannian structure on $\mathcal{M}$ is of step $\gamma+1=2$, and under H1H1-glob, it is of step $\gamma+1\geq 2$. This will always be implied in our statements when we invoke H1H1-loc or H1H1-glob.
We fix in system (5) the control zone $\omega$ satisfying one of the conditions outlined in H1. The singular set $\mathcal{Z}$ satisfies the assumptions of H0, and, without loss of generalities, it is diffeomorphic to $\Omega_{y}=(0,\pi),\ \mathbb{S}^{1}$, or $\mathbb{R}$. The distance $\delta$ to the singularity satisfies H0’.
We recall that we denote by $\mathcal{U}$ a tubular neighborhood constructed as in Section 1.1, under H1, and such that H2 and H3 hold. That is, $\mathcal{U}$ is contained in $\mathbb{M}\setminus\omega$, and is diffeomorphic to $(-L,L)\times\Omega_{y}$ for some $0<L\leq\min(L_{0},R)<\min(d_{\operatorname{sR}}(\omega,\mathcal{Z}),\operatorname{inj}(\mathcal{Z}))$, where $L_{0}>0$ and $R>0$ are respectively such that H2, H3 and H0’ hold.
We also recall that in $\mathcal{U}$, the sub-Riemannian structure is generated by two vector fields $X=\partial_{x}$ and $Y=q(x)r(y)\partial_{y}$, with $q$ and $r$ satisfying the conditions outlined in H2.
To proceed to the proof, we need to characterize how the boundary of such a neighborhood can intersect the boundary of $\mathbb{M}$ under H1H1-glob. In the Lemma below and what follows, the notation $\simeq$ signifies diffeomorphic to.
Lemma 5.1.
Let $\mathcal{U}$ be a tubular neighborhood as prescribed above under H1H1-glob, that is diffeomorphic to $(-L,L)\times\Omega_{y}$ for some $0<L\leq\min(L_{0},R)<\min(d_{\operatorname{sR}}(\omega,\mathcal{Z}),\operatorname{inj}(\mathcal{Z}))$. Then, for $L$ small enough,
  1. (i)
    if $\mathcal{Z}\simeq\mathbb{S}^{1}$ or $\mathbb{R}$, we have $\partial\mathcal{U}\cap\partial\mathcal{M}=\emptyset$,
  2. (ii)
    if $\mathcal{Z}\simeq(0,\pi)$, we have $\partial\mathcal{U}\cap\partial\mathcal{M}\simeq\left([-L,L]\times\{0\}\right)\cup\left([-L,L]\times\{\pi\}\right)$.
Proof.
We need to show that for $L>0$ sufficiently small, we have $\mathcal{U}=\{p\in\mathcal{M},\delta(p)<L\}$. Indeed, assume that this holds.
If $\mathcal{Z}\simeq\mathbb{S}^{1}$, then $\partial\mathcal{U}=\{p\in\mathcal{M},\delta(p)=L\}\simeq\left(\{-L\}\times\mathbb{S}^{1}\right)\cup\left(\{L\}\times\mathbb{S}^{1}\right)$. In this case, since $\mathcal{Z}\subset\operatorname{Int}(\mathcal{M})$, the lemma directly follows as $L$ can be chosen small enough so that $\mathcal{U}\subset\operatorname{Int}(\mathcal{M})$. The same idea goes for $\mathcal{Z}\simeq\mathbb{R}$.
If $\mathcal{Z}\simeq(0,\pi)$, let $p=(0,0)\in\partial\mathcal{U}\cap\partial\mathcal{Z}\cap\partial\mathcal{M}$ by H0, and consider, in coordinates, the curve $c:t\in[0,L]\mapsto c(t)=(0,t)$ which belongs to $\partial\mathcal{U}$. Assume that $c$ has exited $\partial\mathcal{M}$ at some time $t_{0}\in(0,L)$. Then, there exists an open ball around $c(t_{0})$ of radius $\varepsilon>0$ sufficiently small such that $B(c(t_{0}),\varepsilon)\subset\operatorname{Int}(\mathcal{M})$ and $t_{0}+\varepsilon<L$, and a point $p^{\prime}\in B(c(t_{0}),\varepsilon)\setminus\mathcal{U}$. By continuity of $\delta$ from H0’, we have $\delta(p^{\prime})\leq t_{0}+\varepsilon<L$. Hence $p^{\prime}\in\{p\in\mathcal{M},\delta(p)<L\}$ which is a contradiction. The other parts of $\partial\mathcal{U}$ can be treated the same way. Taking $L$ small enough directly concludes the proof.
It remains to prove that for $L\in(0,\min(d_{\operatorname{sR}}(\omega,\mathcal{Z}),\operatorname{inj}(\mathcal{Z}))$ we have $\mathcal{U}=\{p\in\mathcal{M},\delta(p)<L\}$.
The set equality is proved in the case $\mathcal{Z}\simeq\mathbb{S}^{1}$ in [Franceschi2020-gs, Proposition 3.1], [rossi2022relative, Theorem 3.7]. When $\mathcal{Z}\simeq\mathbb{R}$, it still holds thanks to the injectivity radius assumption in H0 (see [Franceschi2020-gs, footnote p.98]).
Let us focus on the remaining case. We obviously have the inclusion $\mathcal{U}\subset\{p\in\mathcal{M},\delta(p)<L\}$, so let us prove the reciprocal one. Let $p_{0}\in\{q\in\mathcal{M},\delta(q)<L\}\setminus\mathcal{U}$. Let $q\in\mathcal{Z}$ be the unique point in $\mathcal{Z}$ such that $\delta(p_{0})=d_{\operatorname{sR}}(p_{0},q)$. Then, there exists $c_{0}:[0,1]\rightarrow\mathcal{M}$ a minimizing geodesic such that $c_{0}(0)=q$ and $c_{0}(1)=p_{0}$. Assume that there exists $p_{1}\in\mathcal{U}$ in the same connected component of $\{q\in\mathcal{M},\delta(q)<L\}$ as $p_{0}$ such that $\delta(p_{1})=d_{\operatorname{sR}}(p_{0},q)=\delta(p_{0})$. We know that $p_{1}$ exists since it is either $p_{1}=(q,\delta(p_{0}))$ or $p_{1}=(q,-\delta(p_{0}))$ depending on where $p_{0}$ is situated. There exists a minimizing normal geodesic $c_{1}$ such that $c_{1}(0)=q$ and $c_{1}(1)=p_{1}$. Observe that the minimizing geodesics $c_{i}$, $i=0,1$, must satisfy $c_{i}^{\prime}(t)=\alpha(t)\nabla\delta(c_{i}(t))$, where $\alpha:[0,1]\rightarrow\mathbb{R}^{+}$ is smooth. Since we assumed smoothness of $\delta$ in H0’, we must have that there exists $\beta>0$ such that $c_{0}^{\prime}(0)=\beta c_{1}^{\prime}(0)$. Since $c_{0}$ and $c_{1}$ are both integral curves of $\nabla\delta$ starting from the same point, we must have that there exists a smooth function $\beta:[0,1]\rightarrow\mathbb{R}^{+}$ such that $\beta(0)=\beta$ and $c_{1}^{\prime}(t)=\beta(t)c_{0}^{\prime}(t)$. That is, $c_{1}$ is a reparametrization of $c_{0}$, and hence $c_{0}(1)=c_{1}(1)$. Thus, $p_{0}=p_{1}$ and the inclusion $\{p\in\mathcal{M},\delta(p)<L\}\subset\mathcal{U}$ follows. This concludes the proof. ∎
Remark 5.2.
What the above proof says, is that assuming smoothness of $\delta$ as in H0’, we only have a unique way of exiting $\mathcal{Z}$ in each side of it, by means of a minimizing geodesic (minimizing the distance to $\mathcal{Z}$), which is a normal one, from any point of the singularity. This geodesic must be an integral curve for $\nabla\delta$. If we exit from $q\in\operatorname{Int}(\mathcal{Z})$, this geodesic exists, is unique, and must be normal (see [Franceschi2020-gs, Proposition 2.7] or [bossio2024tubes, Proposition 2.2]). However, this is not true in all generalities from $q\in\partial\mathcal{Z}$, as it may be done by means of different normal geodesics, or abnormal ones. However, the existence of a normal geodesic is ensured. However, the smoothness assumption from H0’ forces all possibilities of leaving $\mathcal{Z}$ from $q$ to be along the same path. That is, even if $q\in\partial\mathcal{Z}$, this geodesic must be unique (modulo reparametrization) and normal. This in turns forces $\mathcal{U}=\{p\in\mathcal{M},\delta(p)<L\}$.
5.1 Reduction process
We fix from now on $\mathcal{U}$ to be a tubular neighborhood as prescribed by Lemma 5.1 under H1H1-glob, which we recall is diffeomorphic to $(-L,L)\times\Omega_{y}$. Under H1H1-loc, we also fix $\mathcal{U}$ similarly to be a tubular neighborhood of a point $p\in\mathcal{Z}$ such that $\partial\mathcal{U}\subset\operatorname{Int}(\mathcal{M})$, which is always possible. Under H1H1-loc, $\mathcal{U}\simeq\Omega=(-L,L)\times(0,\pi)$ and $\partial\mathcal{U}\simeq\partial\Omega$.
We show in this subsection that null-controllability of system (5) implies the internal null-controllability of initial states in the form domain $D\left(\Delta_{\mathcal{U}}^{1/2}\right)$ of system
$\left\{\begin{array}[]{lcll}\partial_{t}f-\Delta_{\mathcal{U}}f&=&\mathbf{1}_{\omega^{\varepsilon}}(x,y)u(t,x,y),&(t,x,y)\in(0,T)\times\mathcal{U},\\ f(t,x,y)&=&0,&(t,x,y)\in(0,T)\times\partial\mathcal{U},\\ f(0,x,y)&=&f_{0}(x,y),&(x,y)\in\mathcal{U}.\end{array}\right$. (131)
For simplicity of the presentation, we shall denote the elements of $\mathcal{U}$ by $(x,y)$ with $x\in(-L,L)$ and $y\in\Omega_{y}$.
Proposition 5.3.
Under the fixed setting previously prescribed, assume that system (5) is null-controllable in time $T>0$ from $\omega$ that satisfies $H1$-loc. Let $\varepsilon>0$, and define
$\omega^{\varepsilon}=\left\{(x,y)\in\mathcal{U},\quad x\in(-L,-L+\varepsilon)\cup(L-\varepsilon,L),\quad y\in(0,\varepsilon)\cup(\pi-\varepsilon,\pi)\right\}.$
Then, for every $\varepsilon>0$, for every $f_{0}\in D\left(\Delta_{\mathcal{U}}^{1/2}\right)$, there exists a control $u\in L^{2}((0,T)\times\omega^{\varepsilon})$ such that the associated solution of system (131) satisfies $f(T)=0$ in $L^{2}(\mathcal{U})$.
Moreover, if the control zone $\omega$ in system (5) satisfies $H1$-glob, then the same conclusion holds with
$\omega^{\varepsilon}=\left\{(x,y)\in\mathcal{U},\quad x\in(-L,-L+\varepsilon)\cup(L-\varepsilon,L)\right\},$
for every $\varepsilon\in(0,L)$.
We recall that when working under H1H1-loc, our structure is of step $\gamma+1=2$, and under H1H1-glob, it is of step $\gamma+1\geq 2$. In the second part of the above Proposition, if $\gamma+1>2$, and $\varepsilon$ is such that $\omega^{\varepsilon}=\mathcal{U}$, the conclusion of our proof of Proposition 5.3 is no longer true. Using our approach via cutoff arguments, when the control zone $\omega^{\varepsilon}$ intersects the singularity, one needs to be careful on the regularity of the initial states in system (131) that can be steered to 0 by means of $L^{2}$-controls induced by the assumption that system (5) is null-controllable. For a structure of step $k$, one would at least be required to consider initial states in $D\left(\Delta_{\mathcal{U}}^{\frac{k-1}{2}}\right)$. However, in the proof below, some extension arguments would need to be carefully checked. Thus, under H1H1-glob, for simplicity, we restrict $\varepsilon>0$ to be small so that $\omega^{\varepsilon}$ is located within the elliptic region, and so that our argument works when $\gamma+1>2$.
Proof.
Let $T>0$. Assume that we are under H1H1-loc, and that system (5) is null-controllable in time $T>0$ from $\omega$. The following arguments hold due to the analyticity of the semigroup associated to our operators $\Delta$ and $\Delta_{\mathcal{U}}$.
Let $f_{0}\in D\left(\Delta_{\mathcal{U}}^{1/2}\right)$. Extend it by 0 outside of $\Omega$. Then, the function $\mathbf{1}_{\mathcal{U}}f_{0}$ belongs to the form domain $D\left(\Delta^{1/2}\right)$. Indeed, this can be seen by approximating $f_{0}$ by smooth compactly supported functions in $\mathcal{U}$, that are extended outside of $\mathcal{U}$ by zero. Then, as $D\left(\Delta^{1/2}\right)$ is the completion of $C_{c}^{\infty}(\mathcal{M})$ with respect to the horizontal Sobolev norm
$\|u\|_{D(\Delta^{1/2})}^{2}=\int_{\mathcal{M}}\sum_{i=1}^{m}|X_{i}u(p)|^{2}\ d\mu(p),$
where the family $\left\{X_{1},...X_{m}\right\}$ is a generating frame for the sub-Riemannian structure, this concludes.
Now, by assumption, there exists a control $v\in L^{2}((0,T)\times\omega,dtd\mu)$ such that the solution $f_{\operatorname{man}}$ of system (5) with initial state $\mathbf{1}_{\mathcal{U}}f_{0}$ satisfies $f_{\operatorname{man}}(T)=0$. Denote now by $f_{\operatorname{adj}}$ the solution of system (131) with $u=0$ and with initial state $f_{0}$. We set
$f_{\operatorname{tub}}:=\eta\theta f_{\operatorname{adj}}+(1-\theta)f_{\operatorname{man}},$
where $\eta\in C^{\infty}([0,T])$ and $\theta=\theta_{1}\theta_{2}\in C^{\infty}(\mathcal{U})$ are defined by
$\left\{\begin{array}[]{lllr}\theta_{1}(x)&=&1,&x\in[-L,-L+\frac{\varepsilon}{2}]\cup[L-\frac{\varepsilon}{2},L],\\ \theta_{1}(x)&=&0,&x\in[-L+\varepsilon,L-\varepsilon],\\ \theta_{1}(x)&\in&[0,1],&x\in[-L,L],\end{array}\right.\quad\left\{\begin{array}[]{lllr}\theta_{2}(x)&=&1,&x\in[0,\frac{\varepsilon}{2}]\cup[\pi-\frac{\varepsilon}{2},\pi],\\ \theta_{2}(x)&=&0,&x\in[\varepsilon,\pi-\varepsilon],\\ \theta_{2}(x)&\in&[0,1],&x\in[0,\pi],\end{array}\right$.
and
$\left\{\begin{array}[]{lcll}\eta(t)&=&1,&t\in[0,\frac{T}{3}],\\ \eta(t)&=&0,&t\in[\frac{2T}{3},T],\\ \eta(t)&\in&[0,1],&t\in[0,T].\end{array}\right$. (135)
We observe that in $\mathcal{U}$, $f_{\operatorname{tub}}(0)=f_{0}$, $f_{\operatorname{tub}}(T)=0$, and $f_{\operatorname{tub}}$ values zero at the boundary. Moreover, computations in coordinates using (1) show that applying the parabolic operator $\partial_{t}-\Delta_{\mathcal{U}}$ to $f_{\operatorname{tub}}$ produces a source term $h$ localized in $\omega^{\varepsilon}$, that will depend on $f_{\operatorname{adj}}$, $f_{\operatorname{man}}$ and their first derivatives with respect to $x$ and $y$.
To conclude on the proof, we need to make sure that $h\in L^{2}((0,T)\times\omega^{\varepsilon})$. A sufficient condition for this is that $f_{\operatorname{adj}}$ and $f_{\operatorname{man}}$ both belong to $L^{2}\left((0,T);H^{1}(\mathcal{U})\right)$.
Let us first observe that since $f_{0}\in D\left(\Delta_{\mathcal{U}}^{1/2}\right)$, then $f_{\operatorname{adj}}\in L^{2}((0,T),D\left(\Delta_{\mathcal{U}}\right))$. Subelliptic estimates (see e.g. [Laurent2022-ca, Theorem 1.5 eq. (1.6)]), stating that $D(\Delta_{\mathcal{U}})\subset H^{2/k}(\mathcal{U})$ where $k$ is the step of the structure, then infer that for a.e. $t\in[0,T]$, $f_{\operatorname{adj}}(t)\in H^{1}(\mathcal{U})$, since, as we work under H1H1-loc, the step of the structure is $2$. Thus, we indeed have $f_{\operatorname{adj}}\in L^{2}\left((0,T);H^{1}(\mathcal{U})\right)$.
Now, to treat $f_{\operatorname{man}}$, we write it using the Duhamel formula
$f_{\operatorname{man}}(t)=e^{-t\Delta}\mathbf{1}_{\mathcal{U}}f_{0}+\int_{0}^{T}e^{-(t-s)\Delta}\mathbf{1}_{\omega}v(s)\ ds.$
We have that $e^{-\cdot\Delta}\mathbf{1}_{\mathcal{U}}f_{0}\in L^{2}((0,T),D\left(\Delta\right))$, and thus belongs to $L^{2}\left((0,T);H^{1}(\mathcal{U})\right)$ using the subelliptic estimate [Laurent2022-ca, Theorem 1.5 eq. (1.6)] as above. For the non-homogeneous part, that we denote by $F_{\operatorname{man}}$, we use maximal $L^{2}$-regularity of our operator $\Delta$, which gives us the existence of a constant $C>0$ such that
$\|\Delta F_{\operatorname{man}}\|_{L^{2}\left((0,T),L^{2}(\mathcal{M},d\mu)\right)}^{2}\leq C\|\mathbf{1}_{\omega}v\|_{L^{2}\left((0,T),L^{2}(\mathcal{M},d\mu)\right)}^{2}<\infty.$
That is, $F_{\operatorname{man}}\in L^{2}\left((0,T);D(\Delta)\right)$. Thus, using once again the subelliptic estimate [Laurent2022-ca, Theorem 1.5 eq. (1.6)], $F_{\operatorname{man}}\in L^{2}\left((0,T);H^{1}(\mathcal{M})\right)\subset L^{2}\left((0,T);H^{1}(\mathcal{U})\right)$. This concludes on the fact that $f_{\operatorname{man}}\in L^{2}\left((0,T);H^{1}(\mathcal{U})\right)$.
The $L^{2}$-control $u$ in system (131) that steers $f_{0}$ to 0 in time $T$ is therefore given by $h$, and the associated solution is $f_{\operatorname{tub}}$.
Now, under H1H1-glob, we follow the same proof as above until the use of the cutoffs, at which point we simply take $\theta=\theta_{1}$. The resulting non-homogeneous term $h$ acting like the control is now located, for $\varepsilon\in(0,L)$, within the elliptic region. Thus, by local regularity of the solutions in the elliptic region, we indeed have $h\in L^{2}((0,T)\times\omega^{\varepsilon})$. Finally, thanks to Lemma 5.1, the Dirichlet boundary conditions are also satisfied. This fully concludes the proof. ∎
5.2 Proof of Theorem 1.5
We can now proceed to conclude on the proof of Theorem 1.5. Both considerations in the theorem are treated the same way.
Proof of Theorem 1.5.
We illustrate the proof in Figure 3. We assumed that system (5) is null-controllable in time $T>0$ from $\omega$ in the beginning of Section 5.
Let us recall $\mathcal{U}$ is fixed at the beginning of Section 5. Under both considerations outlined in H1, $\mathcal{U}$ is diffeomorphic to, via lets say $\Phi$, $\Omega=(-L,L)\times\Omega_{y}$ with $L>0$ given by Lemma 5.1.
Moreover, without loss of generality, we assumed that
[…]
The diffeomorphism $\Phi:\mathcal{U}\rightarrow\Omega$ induces a unitary transformation
$\begin{array}[]{lccc}S:&L^{2}(\mathcal{U},\mu)&\rightarrow&L^{2}(\Omega,h(x)dxdy)\\ &f&\mapsto&Tf=f\circ\Phi,\end{array}$
and we have
$G:=S\Delta_{\mathcal{U}}S^{-1}=\frac{1}{h(x)}\partial_{x}(h(x)\partial_{x})+q(x)^{2}\partial_{y}(r(y)^{2}\partial_{y}).$ (136)
Thus, by Proposition 5.3, and the above unitary transformation, we have internal null-controllability in time $T$ of initial states $f_{0}\in D\left(G^{1/2}\right)$ of system
$\left\{\begin{array}[]{lcll}\partial_{t}f-(\frac{1}{h(x)}\partial_{x}(h(x)\partial_{x})+q(x)^{2}\partial_{y}(r(y)^{2}\partial_{y}))f&=&u(t,x,y)\mathbf{1}_{\omega^{\varepsilon}}(x,y),&(t,x,y)\in(0,T)\times\Omega,\\ f(t,p)&=&0,&(t,p)\in(0,T)\times\partial\Omega,\\ f(0,p)&=&f_{0}(p),&p\in\Omega,\end{array}\right$. (140)
posed in $L^{2}(\Omega,h(x)dxdy)$, and with the $L^{2}$-control $u$ supported on $\omega^{\varepsilon}$ defined by, for $\varepsilon>0$ arbitrary small,
  1. (i)
    $\omega^{\varepsilon}=\left\{(x,y)\in\Omega,\ x\in(-L,-L+\varepsilon)\cup(L-\varepsilon,L)\right\}\cup\left\{(x,y)\in\Omega,y\in(0,\varepsilon)\cup(\pi-\varepsilon,\pi)\right\}$, under H1H1-loc,
  2. (ii)
    $\left\{(x,y)\in\Omega,\ x\in(-L,-L+\varepsilon)\cup(L-\varepsilon,L)\right\}$, under H1H1-glob.
Consider now the unitary transformation
$\begin{array}[]{lccl}T:&L^{2}(h(x)dxdy)&\longrightarrow&L^{2}(dxdy)\\ &f&\longmapsto&\sqrt{h}f.\end{array}$ (143)
Then, setting $G_{V}:=TGT^{-1}=\partial_{x}^{2}f+q(x)^{2}\partial_{y}(r(y)^{2}\partial_{y}f)+V(x)f$, the observability inequality associated to the problem of internal null-controllability in time $T>0$ of initial states in $D\left(G^{1/2}\right)$ of system (140), is equivalent to the observability inequality (29), associated to the problem of internal null-controllability in time $T>0$ of initial states in $D\left(G_{V}^{1/2}\right)$ of system (19), with $V(x)=\frac{\partial_{x}^{2}(\sqrt{h})}{\sqrt{h}}$.
Therefore, the non null-controllability results presented in Theorem 1.5 are inherited from both Theorems 2.1 and 2.2.
  1. 1.
    Under H1H1-loc, and $\gamma=1$, from Theorem 2.1 we have that for every $\varepsilon>0$ arbitrary small, for any time
    $T<\frac{1}{q^{\prime}(0)}\min\{d_{\operatorname{agm}}(-L+\varepsilon),d_{\operatorname{agm}}(L-\varepsilon)\},$
    there exists $f_{0}\in D\left(G_{V}^{1/2}\right)$ such that the associated solution of system (19) cannot be steered to 0 in time $T$. As $\varepsilon$ can be taken arbitrary small, we get for system (5) that
    $T(\omega)\geq\frac{1}{q^{\prime}(0)}\min\{d_{\operatorname{agm}}(-L),d_{\operatorname{agm}}(L)\}.$
  2. 2.
    Under $H1$-glob, by Theorem 2.2, when $\gamma>1$ we obviously never have null-controllability, since system (19) is never null-controllable, and for $\gamma=1$, we obtain the same lower bound for $T(\omega)$ as above.
(a) The reduction process under H1H1-glob.
(b) The reduction process under H1H1-loc.
Figure 3: The reduction process under H1. The singularity is in red, the control zones in olive, and the tubular neighborhood $\mathcal{U}$ in light blue. The blue paths correspond to $\partial\mathcal{U}\setminus\partial\mathbb{M}$. The dashed part of $\mathcal{U}$ correspond to where the control is supported after applying Proposition 5.3. The arrows represent the passage into coordinates representation, with in olive on the tensorized domains the support of the control for the problem of internal controllability, which is the image by $\Phi$ of the dashed parts.
6 Applications, comments, and open problems
The Grushin operator without warped product assumption. Consider $\mathcal{M}=(-1,1)\times(0,\pi)$ with the Lebesgue measure. Set $X=\partial_{x}$, $Y=\tilde{q}(x,y)\partial$. The sub-Laplacian writes $\Delta=\partial_{x}^{2}+\partial_{y}(\tilde{q}(x,y)^{2}\partial_{y})$. Although there is no result for the heat equation associated to this operator, Theorem 1.5 says that to obtain a negative result, it is sufficient that $\tilde{q}$ locally writes as $\tilde{q}(x,y)=q(x)r(y)$. Having a result for $\Delta$, even in the particular case of the rectangle, is an open problem. The usual Fourier techniques are not valid in this case.
The Grushin sphere. We retrieve with the strategy of Theorem 1.5 the negative result of Tamekue [tamekue2022null]. Namely, consider $\mathcal{M}=\mathbb{S}^{2}$, endowed with the sub-Riemannian structure generated by the vector fields $X=-z\partial_{y}+y\partial_{z}$, $Y=-z\partial_{x}+x\partial_{z}$. Observe that they are linearly independent outside $\{z=0\}$, but $\{X,Y\}$ is bracket-generating since $X,Y]=-y\partial_{x}+x\partial_{y}$. Moreover, H0 is satisfied. We also endow $\mathbb{S}^{2}$ with the restricted Lebesgue measure. Consider system (5) with $\omega$ consisting of two symmetric, with respect to $\{z=0\}$, horizontal crowns as in [tamekue2022null, Fig. 1]. For example $\omega=\{(x,y,z)\in\mathbb{S}^{2},|z|>a\}$ for some $a>0$. Then, reducing the problem in $\mathbb{S}^{2}\setminus\omega$, in spherical coordinates, the restricition of the Lebesgue measure writes $\cos(x)dxdy$, and we have that $\Delta=\frac{1}{\cos(x)}\partial_{x}(\cos(x)\partial_{x})+\tan(x)^{2}\partial_{y}^{2}$ on $(-a,a)\times\mathbb{S}^{1}$. Then, by parity of the function $\tan$, our result shows that $T(\omega)\geq\frac{1}{q^{\prime}(0)}\int_{0}^{a}\tan(s)\ ds=\ln\left(\frac{1}{\cos(a)}\right)$, which coincides with the negative result in [tamekue2022null]. This strategy would also hold if the two components of the control zone are not at the same nonnegative distance of the singularity.
The Grushin equation on the real line and lack of optimality. Consider the Grushin equation on $\mathcal{M}=\mathbb{R}\times(0,\pi)$, with the Lebesgue measure, controlled on vertical strips. Theorem 1.5 says that we can reduce the analysis of the negative results to the case $(-L,L)\times(0,\pi)$. We can emphasize that when $\lim_{|x|\rightarrow+\infty}|q(x)|=+\infty$, and $q$ satisfies H2, the strategy proposed in the case where $\Omega_{x}$ is bounded extends to $\mathbb{R}$, and also if we perturb the equation by a potential $V\in L_{loc}^{\infty}(\mathbb{R})$ bounded below. The strategy of Theorem 2.2 extended to $\mathbb{R}$ and Theorem 1.5 both give the same results. Nonetheless, it is important to notice that in both cases we obtain a lower bound for the minimal time of null-controllability, even when the equation is never null-controllable (for example when we do not control in one of the connected component of $\mathcal{M}\setminus\mathcal{Z}$).
Non-symmetry of the tubular neighborhood. While Theorems 2.1 and 2.2 assume that $\Omega_{x}$ is symmetric with respect to zero, they remain valid when substituting (L,L)(-L,L) with any other interval containing 0 in its interior, up to some minor adjustments, in particular for Theorem 2.2. For Theorem 2.1, the result comes from [darde2023null] where the symmetry is never assumed. For Theorem 2.2, the argument for exponential decay of the classical eigenfunctions in Proposition 3.6 is no longer valid on the whole interval. That is, by lack of parity, Remark 3.8 does not hold. One can either reapply the proof to negative $x$, or obtain an exponential decay argument as in Proposition 4.1. In any case, the asymptotic on the eigenvalues still hold, and so does Theorem 2.2. As a consequence, we can choose $\mathcal{U}$ to be diffeomorphic to $(-L_{-},L_{+})\times\Omega_{y}$.
Positive results on $\mathcal{M}$. Although we only focus on negative results, positive results are achievable in our settings when $\gamma=1$. Indeed, if a positive result is achieved in $\mathcal{U}$ in time $T>0$ from $\omega$, it is possible to extend it to the whole manifold through cutoffs arguments. However, the resulting control zone $\omega^{\prime}\subset\mathcal{M}$ for the system posed on $\mathcal{M}$ will have to be taken accordingly to the achievable positive results for the system posed on $\mathcal{M}\setminus\mathcal{U}$. For instance, under H1H1-loc, to the best of our knowledge, we can only state that the heat system posed on $\mathcal{M}\setminus\mathcal{U}$ is null-controllable in any time $T>0$ from $\mathcal{M}\setminus\mathcal{U}$ by dissipativity. Under H1H1-glob however, we can take advantage of the uniform parabolicity of the heat outside $\mathcal{U}$.
Nevertheless, to obtain null-controllability within $\mathcal{U}$, we must also require that $\inf_{\mathcal{U}}q^{\prime}(x)>0$, which is a stronger assumption than H2 (see [beauchard2020minimal]). Proving positive results for the Grushin equation on rectangular domains without the monotony assumption on $q$ is still an open problem when the control zone does not act on both side of the singularity.
Complement of a horizontal strip. The analogue on manifolds of controlling in the complement of a horizontal strip on a rectangular domain, is to say that there exists a normal geodesic to the singularity that never intersects the control zone. Although we expect to never have null-controllability of any $\gamma\geq 1$, Theorem 1.5 does not cover this case correctly. It only gives a lower bound for $T(\omega)$. It is also important to say that a negative result when $\gamma>1$ in this setting on the rectangle has never been obtained, and is an open problem. If such a result is obtained, Theorem 1.5(ii) would hold under H1H1-loc.
Acknowledgment
The author would like to express his gratitude to his PhD advisors, Pierre Lissy and Dario Prandi, for their guidance, insightful discussions, and feedbacks throughout this work. The author also wishes to extend his thanks to Armand Koenig for the discussions regarding his series of papers on the non-observability of the Grushin equation, as well as for the exchanges concerning their generalization, and to anonymous reviewers whose comments greatly contributed to improving a previous version of this paper.